### v1

# Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits

Abstract
This paper is motivated by recent developments in the linear bandit literature, which have revealed a discrepancy between the promising empirical performance of algorithms such as Thompson sampling and Greedy, when compared to their pessimistic theoretical regret bounds. The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances. To address this, we propose a new data-driven technique that tracks the geometry of the uncertainty ellipsoid, enabling us to establish an instance-dependent frequentist regret bound for a broad class of algorithms, including Greedy, OFUL, and Thompson sampling. This result empowers us to identify and “course-correct” instances in which the base algorithms perform poorly. The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$, while retaining most of the desirable properties of the base algorithms. We present simulation results to validate our findings and compare the performance of our algorithms with the baselines.
1 Introduction
Multi-armed bandits (MABs) have garnered significant attention as they provide a well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems. In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards. This paper specifically focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases. In this model, the system parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield rewards equal to the inner product of $\theta^{\star}$ and the chosen action, along with mean-zero, subgaussian noise. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total rewards achieved by the policy and the maximum achievable reward. Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985). Other popular strategies include Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
In the linear bandit setting, two types of regret are considered. The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al. (2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to logarithmic factors. On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of noise and algorithm.
In the frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al. (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound. However, for a frequentist variant of LinTS referred to as TS-Freq, Agrawal and Goyal (2013) and Abeille et al. (2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret by inflating the posterior distribution. This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$. Recent work by Hamidi and Bayati (2020a) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved. While no general theoretical guarantee exists for the Greedy algorithm in linear bandit problems, it is known that both Greedy and LinTS perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
Despite the theoretical gaps in performance, LinTS has shown strong empirical performance, indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently. Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021). Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al. (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency. This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify problematic instances where LinTS and Greedy could potentially fail, and apply a “course-correction” to ensure competitive frequentist regret bounds? And can this be achieved without compromising their impressive empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
  • Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid. This technique is essential in maximizing the utilization of historical data, thereby enabling the “course-correction”.
  • We derive a general instance-dependent frequentist regret bound for the proposed POFUL algorithms, a family of algorithms including OFUL, LinTS, TS-Freq, and Greedy as special cases. The bound can be efficiently calculated in a data-driven manner based on the observation history.
  • We propose variants of LinTS and Greedy that enjoy the minimax optimal frequentist regret, while retaining most of the desirable properties such as computational efficiency.
The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid. This is in stark contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their overly narrow concentration on comparatively rudimentary summary statistics.
Related literature.
The ideas presented in this paper bear similarities to existing literature that explores different forms of exploration (Russo and Van Roy, 2016; Kirschner and Krause, 2018) or recent works by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b) that aim to utilize observed data for reducing exploration. However, the methodology and the manner in which the data is utilized in this paper differ significantly. For instance, Bastani et al. (2021) examines the minimum eigenvalue of the sample covariance matrix which is a single-parameter summary statistic of the observed data, while Hamidi and Bayati (2020b) incorporates information from the (one-dimensional) confidence intervals around the reward values. In contrast, our approach leverages the complete geometric information of the $d$-dimensional confidence ellipsoid.
2 Setup and Preliminaries
Notations.
[…]
Problem Formulation.
We consider the stochastic linear bandit problem. Let $\theta^{\star}\in\mathbb{R}^{d}$ be a fixed but unknown parameter. At each time $t\in[T]$, a policy $\pi$ selects action $x_{t}$ from a set of action $\mathcal{X}_{t}\subset\mathbb{R}^{d}$ according to the past observations and receives a reward $r_{t}=\langle{x_{t}},{\theta^{\star}}\rangle+\varepsilon_{t}$, where $\varepsilon_{t}$ is mean-zero noise with a distribution specified in Assumption 3 below. We measure the performance of $\pi$ with the cumulative expected regret $\mathcal{R}(T)=\sum_{t=1}^{T}\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x_{t}},{\theta^{\star}}\rangle,$ where $x^{\star}_{t}$ is the best action at time $t$, i.e., $x^{\star}_{t}=\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$. Let $\mathcal{F}_{t}$ be a $\sigma$-algebra generated by the history $(x_{1},r_{1},\ldots,x_{t},r_{t})$ and the prior knowledge, $\mathcal{F}_{0}$. Therefore, $\{\mathcal{F}_{t}\}_{t\geq 0}$ forms a filteration such that each $\mathcal{F}_{t}$ encodes all the information up to the end of period $t$.
2.1 Assumptions
We make the following assumptions that are standard in the relevant literature.
Assumption 1 (Bounded parameter)
The unknown parameter $\theta^{\star}$ is bounded as $\|\theta^{\star}\|\leq S$, where $S>0$ is known.
Assumption 2 (Bounded action sets)
The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$. Without loss of generality, we normalize the actions such that $\mathcal{X}_{t}\subset{\mathcal{S}}_{d-1}$ for all $t\in[T]$.
Assumption 3 (Noise)
The noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ is conditionally mean-zero and $R$-subgaussian, where $R$ is known. Formally, for all real valued $\lambda$, $\mathbb{E}\left[e^{\lambda\varepsilon_{t}}|\mathcal{F}_{t}\right]\leq\exp\left(\lambda^{2}R^{2}/2\right)$. This condition implies that $\mathbb{E}\left[\varepsilon_{t}|\mathcal{F}_{t}\right]=0$ for all $t\geq 1$.
2.2 Regularized Least Square and Confidence Ellipsoid
Here, we review useful frequentist tools developed in Abbasi-Yadkori et al. (2011).
Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards. In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator. Fix a regularization parameter $\lambda$, the design matrix, and the Regularized Least Square (RLS) estimate are respectively given by
$V_{t}=\lambda I+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
The following proposition in Abbasi-Yadkori et al. (2011) states that the RLS estimate $\widehat{\theta}_{t}$ concentrates around $\theta^{\star}$ properly.
Proposition 1 (Theorem 2 in Abbasi-Yadkori et al. (2011))
Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that
$\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta),\quad\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$
for all $x\in\mathbb{R}^{d}$, where
$\beta_{t}^{RLS}(\delta)=R\sqrt{2\log\frac{(\lambda+t)^{d/2}\lambda^{-d/2}}{\delta}}+\sqrt{\lambda}S.$ (2)
Hereafter, we might omit the dependence of $\beta_{t}$ on $\delta$ if there is no ambiguity. Propostion 1 enables us to construct the following sequence of confidence ellipsoids.
Definition 1
Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as $\mathcal{E}_{t}^{RLS}(\delta)=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)\}$.
The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret. Basically, this proposition is saying that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
Proposition 2 (Lemma 11 in Abbasi-Yadkori et al. (2011))
If $\lambda>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda I)}\leq 2d\log\left(1+\frac{t}{\lambda}\right).$
3 POFUL Algorithms
In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
Definition 2
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$,
$\mathbb{P}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]\geq 1-\delta^{\prime},$
where we define the pivot ellipsoid as $\mathcal{E}_{t}^{PVT}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t}(\delta)\right\}$.
At each time $t$, POFUL chooses the action that maximizes the optimistic reward
$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$ (3)
A pseudocode representation of POFUL is provided in Algorithm 1.
Algorithm 1 POFUL
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Sample a feasible pivot $\widetilde{\theta}_{t}$ with respect to $\iota_{t}$ according to Definition 2
     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+\widetilde{x}_{t}\widetilde{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}\widetilde{x}_{s}r_{s}$.
  end for
The intuition behind POFUL is combining the exploration mechanism of OFUL and LinTS. OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores by sampling from the confidence ellipsoid. We allow POFUL to select an arbitrary pivot from $\mathcal{E}_{t}^{PVT}$ to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$. By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
OFUL
For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$ Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
LinTS and TS-Freq
Linear Thompson Sampling (LinTS) algorithm is a generic randomized algorithm that samples from a distribution constructed from the RLS estimate at each step. At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, and $\iota^{TS}_{t}$ is an inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability.
Definition 3
Fix any $\delta\in(0,1)$. $\mathcal{D}^{SA}(\delta)$ is a distribution such that $\mathbb{P}_{\theta\sim\mathcal{D}^{SA}(\delta)}\left[\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq 1\right]\geq 1-\delta.$
LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$. We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm. On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), denoted by TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
Greedy
Greedy is a special case of POFUL where $\iota_{t}=\tau_{t}=0$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$ for all $t\in[T]$.
4 Frequentist Regret Analysis of POFUL
In this section, we present the frequentist regret analysis of POFUL algorithms. We first introduce useful concentration events that hold with high-probability.
Definition 4
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. We define $\beta^{PVT}_{t}(\delta^{\prime})\coloneqq\iota_{t}(\delta)\beta^{RLS}_{t}(\delta^{\prime})$ and
$\ \widehat{\mathcal{A}}_{t}$$\coloneqq\left\{\forall s\leq t:\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\left\{\forall s\leq t:\left\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\right\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})\right\}.$
We also define $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$
Proposition 3
Under Assumptions 1, 2 and 3 , we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta.$
In the following, we decompose the regret of POFUL into three terms, $\mathcal{R}^{PE}(T)$, $\mathcal{R}^{PVT}(T)$ and $\mathcal{R}^{RLS}(T)$, and sketch how they are bounded separately. Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the benchmark action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
$\eqqcolon\mathcal{R}^{PE}(T)+\mathcal{R}^{PVT}(T)+\mathcal{R}^{RLS}(T).$
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$. Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$. Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, later we see this could be problematic when bounding $\mathcal{R}^{PE}(T)$.
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$. For OFUL, this term is bounded properly with high probability by the construction of OFUL actions. For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$. In Abeille et al. (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e. $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$. For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures. The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy.
In fact, non-optimistic samples might be able to control the pessimism term. Our method does not force choosing optimistic samples and hence there is no need to inflate the posterior. As a result, it does not introduce a $\sqrt{d}$-gap in the regret. Instead, we introduce a measurement of the “quality” of POFUL actions with respect to the regret, and develop a computable upper bound for this quality using geometric approaches. The main merits are an instance-dependent regret bound for POFUL that matches the minimax optimal regret in some settings, and variants of LinTS and Greedy with provable frequentist regret bound that is minimax optimal up to only a constant factor.
4.1 An Instance-Dependent Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix D.2.
Proposition 4
Suppose ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
On the RHS, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret. To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$ norm.
Definition 5
For each $t\geq 1$, let $\widetilde{x}_{t}$ and $x_{t}^{\star}$ respectively denote the action chosen by POFUL and the optimal action. We define the sensitivity ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$. We also define the regret ratio at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
Recall the Cauchy–Schwarz inequality $\langle{x},{\theta_{1}-\theta_{2}}\rangle\leq\|x\|_{V_{t}^{-1}}\|\theta_{1}-\theta_{2}\|_{V_{t}}$ , intuitively $\alpha_{t}$ serves as a measure of how sensitive the reward $\langle{\widetilde{x}_{t}},{\theta}\rangle$ is compared with that of the optimal action $\langle{x^{\star}_{t}},{\theta}\rangle$, with respect to the change of the parameter $\theta$ in the $V_{t}$-norm.
The intuition behind the definition for $\mu_{t}$ is connecting POFUL and OFUL. It provides an upper bound for the ratio of instantaneous regrets incurred by POFUL and the OFUL-type upper bound. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$. In this sense, our definition of $\mu_{t}$ and $\alpha_{t}$ can be extended to OFUL. One feasible choice is setting $\mu_{t}\equiv 2$ for OFUL and therefore $\alpha_{t}=(\mu_{t}-1-\iota_{t}-\tau_{t})/(1+\iota_{t}-\tau_{t})=1$ (recall $\iota_{t}=0$ and $\tau_{t}=1$ for OFUL). The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix D.3.
Theorem 1 (Oracle frequentist regret bound for POFUL)
Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a frequentist regret of
$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda}\right)}\beta_{T}^{RLS}(\delta^{\prime}).$ (4)
Remark 1
A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al. (2017), where they indicate their proof of regret bound can be reproduced if the algorithm samples $\widetilde{\theta}_{t}$ that satisfies $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$. They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation. Our work answers this question positively in the next section by investigating the property of (the reciprocal of) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
Remark 2
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy. Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL. This is indeed our aim in the next section.
5 A Data-Driven Approach to Calibrating POFUL
In this section, we present the main contribution of this work which provides a data-driven approach to calibrating POFUL. We’ll focus on LinTS and Greedy as they lack theoretical results of frequentist regret bound. In these settings, we have $\tau_{t}=0$ and hence $\mu_{t}\coloneqq(1+\alpha_{t})(1+\iota_{t})$. Note that $\iota_{t}$ is the inflation parameter that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound for the sensitivity ratio $\alpha_{t}$, which turns into an upper bound for the regret ratio $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
In the following, we first construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$. Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids. Subsequently, we show how these bounds can be utilized to calibrate POFUL. Specifically, we propose the TS-MR and Greedy-MR algorithms that attain a provable minimax-optimal frequentist regret up to constant factors, and yet keep the majority of their original properties such as computational efficiency and low regret in most instances.
5.1 Upper Bounding the Sensitivity Ratio
We consider two scenarios: (1) Continuous action space and (2) Discrete action space. For sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for the first scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). The analysis of the second scenario is deferred to Appendix A.
Continuous action space.
Note that in this case, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$, we only need to study the property of the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$. A direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the ellipsoids. Specifically, we have
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (5)
At first glance, one might think that (5) is too conservative to provide useful calibration for POFUL. However, it turns out that the lower bound is pretty tight in practice. As is illustrated in Figure 1. The intuition is that, when POFUL has implemented sufficient exploration so that the confidence ellipsoid $\mathcal{E}_{t}$ is small enough, the potentially optimal actions concentrate in a small cap $\mathcal{C}_{t}$ on the unit hypersphere ${\mathcal{S}}_{d-1}$. All actions within $\mathcal{C}_{t}$ roughly point to the same direction and thus have similar $V_{t}^{-1}$-norm. The following main theorem of this paper derives a upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Appendix B.
Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$.
The following assumption simplifies the proof, and we remark that it holds almost surely.
Assumption 4
For all $t$, $V_{t}$ has distinct eigenvalues, i.e. $\lambda_{1}(V_{t})>\lambda_{2}(V_{t})\dots>\lambda_{d}(V_{t})\geq 0$.
Theorem 2
Under Assumpsion 4, define $\beta^{PVT}_{t}=\iota_{t}\beta^{RLS}_{t}$ and
[…]
where we let
$\quad m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t}/\lambda_{d}(V_{t}))^{2},\quad M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t})^{2})$
and $k\in[d]$ is such that $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$. Then for all $t\in[T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds for all $s\leq t$ that
$\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s},\quad\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s},\quad\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}.$
To better understand what Theorem 2 implies, we discuss some special cases.
Case 1: a pure exploration regime.
When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$. Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$. Specifically, if $V_{t}=D\mathbb{I}_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
Case 2: linear structures.
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$. To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding sensitivity ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
Proposition 5
Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{\max}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
5.2 TS-MR and Greedy-MR Algorithms
In the following, we propose the algorithm Linear Thompson Sampling with Maximum (Regret) Ratio (TS-MR). The intuition is to detect the performance of LinTS by estimating $\mu_{t}$ and avoid bad LinTS actions by switching to OFUL occasionally.
Specifically, at each time $t$, TS-MR calculates an estimate $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Appendix E.
By construction, TS-MR and Greedy-MR have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
Corollary 1
TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
6 Simulations
(a) example 1
(b) example 2
(c) example 3
Figure 2: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm. At each time $t$, we generate 100 i.i.d. random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$. This is a basic example of standard stochastic linear bandit problems without any extra structure. We set the threshold $\mu=8$ for TS-MR and Greedy-MR. We average simulation results over 20 independent runs. The results are summarized in Figure 1(a). TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$. At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0. Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks. This problem is equivalent to a 10-armed contextual bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. Again, we average simulation results over 20 independent runs. The results are summarized in Figure 1(b). In this setting, Greedy performs suboptimally due to a lack of exploration for all arms. Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects bad regret ratio $\widehat{\mu}_{t}$.
Example 3. Prior mean mismatch (Hamidi and Bayati, 2020a).
This is an example in which LinTS is shown to incur linear Bayesian regret. We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$. It is shown in Hamidi and Bayati (2020a) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$. Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action. We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. We carried out 20 independent runs. The results are summarized in Figure 1(c). We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
7 Conclusion
In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, and Greedy as special cases. Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm. Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest. As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail. We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
Appendix A Discrete Action Space
In this scenario, we consider upper bounding the sensitivity ratio for discrete action space. Compared with the unit hypersphere ${\mathcal{S}}_{d-1}$ considered in Section 5.1 , the main difference is that the oracle optimal action $\theta/\|\theta\|$ for $\theta$ might not lie in the action space $\mathcal{X}_{t}$, so we don’t have a closed form for the attainable optimal action $x^{\star}_{t}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$.
To tackle this problem, we need to characterize the property of all potentially optimal actions. We provide two approaches to solve this problem. This first approach works for the setting where the cardinalities of the action sets are relatively small. The second approach utilizes the statistical properties of the action sets and works for the setting where the cardinalities of the action sets are large.
Small cardinality setting.
Note that the optimistic reward of POFUL when $\tau_{t}=0$ is $\langle{x},{\widetilde{\theta}_{t}}\rangle$, where $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$ is an arbitrary pivot parameter. We have $L_{t}(x)\leq\langle{x},{\widetilde{\theta}_{t}}\rangle\leq U_{t}(x)$, where we define the confidence lower and upper bounds as
$L_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}),\quad U_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$
The potentially optimal $x$ that might optimize $\langle{x},{\widetilde{\theta}_{t}}\rangle$ should satisfy
$U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}\ L_{t}(y),$
i.e. $x$ is not dominated by any other action in $\mathcal{X}_{t}$. Therefore, define
$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\}.$
An upper bound for $\alpha_{t}$ can be calculated as
$\alpha_{t}\leq\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (6)
Large cardinality setting.
In this setting, we characterize the distance between $\theta/\|\theta\|$ and $x^{\star}_{t}(\theta)$. We first introduce the notion of Euclidean covering radius.
Definition 6
The Euclidean covering radius $\rho_{t}$ of a set $\mathcal{X}_{t}\subseteq{\mathcal{S}}_{d-1}$ is given by
$\rho_{t}=\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\coloneqq\max_{y\in{\mathcal{S}}_{d-1}}\min_{x\in\mathcal{X}_{t}}\|y-x\|.$
With the help of the Euclidean covering radius, we prove the following variant of Theorem 2. The proof is postponed to Appendix D
Theorem 3
Under Assumpsion 4, let $\{\rho_{t}\}$ be the sequence of Euclidean covering radius of $\mathcal{X}_{t}$, define $\Psi_{t},\Phi_{t}$ as in Theorem 2. Then for all $t\in[1,T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds that
$\alpha_{s}\leq\frac{\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}{\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}$ (7)
for all $s\leq t$.
To justify this new lower bound, it remains to show that $\rho_{t}$ is small enough under proper conditions. For the sake of a better illustration, we focus on action sets that consist of uniformly and independently distributed points on ${\mathcal{S}}_{d-1}$. The following proposition shows that the Euclidean covering radius has an high-probability upper bound that decays at the order of $\widetilde{\mathcal{O}}(N^{-1/d+o(1)})$, where $N$ is the cardinality of actions.
Proposition 6 (Corollary 2.9 in Reznikov and Saff [2016])
Suppose $\mathcal{X}$ is a set of $N$ uniformly and independently distributed random points on the unit sphere ${\mathcal{S}}_{d-1}$, for any $\kappa>0$, there exist constants $C_{\kappa,d}>0$, such that
$\mathbb{P}\left[\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\leq C_{\kappa,d}\left(\frac{\log N}{N}\right)^{1/d}\right]\geq 1-N^{-\kappa}.$
Appendix B Sketch of the Proof of Theorem 2
In this section, we present the proof of Theorem 2. In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
The proof of Theorem 2 decomposes into two steps. First, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below. Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid. Second, we show that upper and lower bounds for the $V_{t}$-norm can be transferred into upper and lower bounds for the $V_{t}^{-1}$-norm. Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound. We postpone the detailed proof for all lemmas in this section to Appendix C.
Step 1. Bounding the $V_{t}$-norm of actions.
We illustrate our geometric approach in $\mathbb{R}^{2}$ in Figure 3. We first upper bound the $V_{t}$ norm for optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$. In the following, we define $\phi_{t}=\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}-\beta_{t}^{2}}$.
Lemma 1
For any $\theta\in\mathcal{E}_{t}$, we have $\|x^{\star}(\theta)\|_{V_{t}}\leq\|\widehat{\theta}_{t}\|_{V_{t}^{2}}/\phi_{t}$.
The proof is based on the observation that for any $\theta\in\mathcal{E}_{t}$, the line $\{t\theta:t\in\mathbb{R}\}$ (orange dashed lines in Figure 3) and ellipsoid $\mathcal{E}_{t}$ (in green) must have at least one intersection point.
In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$ (blue dashed curve). We can check this intersection is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$. A lower bound for the $V_{t}$-norm is given by the following lemma.
Lemma 2
For any $\theta\in\mathcal{H}_{t}$, we have
$\|x^{\star}(\theta)\|_{V_{t}}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
The proof is based on the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$. Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
Figure 3: Illustration of Step 1 in $\mathbb{R}^{2}$.
Step 2. Bounding the $V_{t}^{-1}$-norm of actions
The following lemma bounds the $V_{t}^{-1}$-norm of an action $x$ given a range of its $V_{t}$-norm.
Lemma 3
Suppose $V$ has distinct eigenvalues $\lambda_{1}(V)>\lambda_{2}(V)\dots>\lambda_{d}(V)$ and $x\in\mathcal{B}_{d}$ satisfies $m\leq\|x\|_{V}^{2}\leq M$, where $M\geq 0$ is such that $\lambda_{k}(V)\leq M<\lambda_{k+1}(V)$. We have
$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$ (8)
The proof is based on writing $x$ as a linear combination of the eigenvectors of $V_{t}$, $x=\sum_{i=1}^{d}w_{i}u_{i}$, and convert the original problem into a linear programming problem with constraints using the expressions $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
Combining the two steps, we complete the proof of Theorem 2.
Appendix C Proof of Lemmas for Theorem 2
C.1 Proof of Lemma 1
Let $\theta=tx$ where $x$ is any unit vector in $\mathbb{R}^{d}$ and $t\in\mathbb{R}$ is a scalar. Consider the equation that characterizes the intersection $\{tx:t\in\mathbb{R}\}\cap\mathcal{E}_{t}$, namely $(tx-\widehat{\theta}_{t})^{\top}V_{t}(tx-\widehat{\theta}_{t})=\beta_{t}^{2}$. Equivalently, we have $t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=0$. This quadratic equation of $t$ has at least one root if the discriminant is non-negative, i.e. $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$. Then by direct computation,
$\|x\|_{V_{t}}\leq\frac{\sqrt{(x^{\top}V_{t}\widehat{\theta}_{t})^{2}}}{\phi_{t}}\leq\frac{\sqrt{x^{\top}x}\sqrt{(\widehat{\theta}_{t})^{\top}V_{t}^{\top}V_{t}\widehat{\theta}_{t}}}{\phi_{t}}=\frac{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}}{\phi_{t}}.$
Note that $\theta\in\mathcal{E}_{t}$ if and only if $\theta=tx$ for some $x$ that satisfies $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, we conclude the proof. $\square$
C.2 Proof of Lemma 2
Note that for any $\theta\in\mathcal{H}_{t}\subset\mathcal{E}_{t}$, it holds that $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$. Also, by the construction of $\mathcal{H}_{t}$, we have $\|\theta\|_{V_{t}}=\phi_{t}$. Then we conclude the proof by the definition
$\|x^{\star}(\theta)\|_{V_{t}}=\frac{\|\theta\|_{V_{t}}}{\|\theta\|}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
$\square$
C.3 Proof of Lemma 3
Let $V=U\Lambda U^{\top}$ be the eigendecomposition of $V$, so that the columns of $U$, $\{u_{1},\dots,u_{d}\}$, form an orthonormal basis of $\mathbb{R}^{d}$ and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$ is a diagonal matrix. We can write $x$ as a linear combination of $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$: $x=\sum_{i=1}^{d}w_{i}u_{i}$, where $\sum_{i=1}^{d}w_{i}^{2}=1$. By direct computation, we have $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
We would like to study the range of $\|x\|_{V^{-1}}^{2}$ when $\|x\|_{V}^{2}$ is bounded as $m\leq\|x\|_{V}^{2}\leq M$. First, let’s focus on maximizing $\|x\|_{V^{-1}}^{2}$, it suffices to solve the LP problem
maximize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}\quad\text{s.t.}\quad\forall i,a_{i}\geq 0,\ \sum_{i=1}^{d}a_{i}=1,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\leq M.$
The Lagrangian is given by
$L(d,\mu,\eta)=\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}+\mu(1-\sum_{i=1}^{d}a_{i})+\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)+\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})+\sum_{i=1}^{d}\kappa_{i}a_{i}.$
The KKT conditions are given by
$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{d}a_{i}=1,\sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{d}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$
Since $\lambda_{i}$’s are distinct, to satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices. Hence, $a_{i}$ can only be non-zero for at most two indices, denoted by $a_{i}$ and $a_{j}$. We write $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$ and hence $\lambda_{i}>\lambda_{j}$. Let $y=\|x\|_{V}^{2}\in[\lambda_{j},\lambda_{i}]$, we have $a\lambda_{i}+(1-a)\lambda_{j}=y$ and $a\lambda_{i}^{-1}+(1-a)\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
By direct computation, $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$, and clearly,
$\begin{cases}\frac{\partial f}{\partial y}&=-\frac{1}{\lambda_{i}\lambda_{j}}<0,\\ \frac{\partial f}{\partial\lambda_{i}}&=(\frac{y}{\lambda_{j}}-1)\frac{1}{\lambda_{i}^{2}}>0,\\ \frac{\partial f}{\partial\lambda_{j}}&=(\frac{y}{\lambda_{i}}-1)\frac{1}{\lambda_{j}^{2}}<0.\end{cases}$
Then the maximum of $\|x\|_{V^{-1}}^{2}$ is obtained by setting $\lambda_{i}=\lambda_{1}$, $\lambda_{j}=\lambda_{d}$, $y=m$, and we have
$x^{*}_{\max}$ $=\frac{m-\lambda_{d}}{\lambda_{1}-\lambda_{d}}\lambda_{1}+\frac{\lambda_{1}-m}{\lambda_{1}-\lambda_{d}}\lambda_{d},\quad\|x^{*}_{\max}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$
Similarly, by investigating the KKT conditions for the LP problem that minimize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}$, the minimum of $\|x\|_{V^{-1}}^{2}$ is obtained by some $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$, $\lambda_{i}>\lambda_{j}$. Setting $\lambda_{u}=\min\{\lambda_{i}\geq M\}$, $\lambda_{l}=\max\{\lambda_{i}\leq M\}$, $y=M$, and we have
$x^{*}_{\min}$ $=\frac{M-\lambda_{l}}{\lambda_{u}-\lambda_{l}}\lambda_{u}+\frac{\lambda_{u}-M}{\lambda_{u}-\lambda_{l}}\lambda_{l},\quad\|x^{*}_{\min}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{u}}+\frac{1}{\lambda_{l}}-\frac{M}{\lambda_{u}\lambda_{l}}.$
This concludes the proof. $\square$
Appendix D Other Proofs
D.1 Proof of Proposition 3
By Propostion 1, we have
${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
$\geq 1-\frac{\delta}{2}.$
Similarly, by Definition 2, we have
${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
$\geq 1-\frac{\delta}{2}.$
Combining the two inequalities above, we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta$$\square$
D.2 Proof of Proposition 4
Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (9)
In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. By direct computation
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $=\left(\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle\right)+\left(\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\right)$
$\quad+\left(\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle\right)+\left(\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right).$
To bound the second term on the right hand side, recall by Equation (3) the POFUL action $\widetilde{x}_{t}$ satisfies
$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (9). As a result, we have
$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}),$
$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}).$
Combing all terms above, we have
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
$\square$
D.3 Proof of Theorem 1
We formally prove Theorem 1 for completeness. The proof techniques are developed in previous papers [Abbasi-Yadkori et al., 2011, Agrawal and Goyal, 2013, Abeille et al., 2017].
Throughout the proof, we condition on the event $\mathcal{A}_{T}$, which holds with probability $1-\delta$ by Proposition 3. Applying Proposition 4, we obtain
$\mathcal{R}(T)$ $\leq\sum_{t=1}^{T}\left(\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right)\mathbb{I}\{\mathcal{A}_{t}\}$
$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
Recall $\|x^{\star}_{t}\|_{V^{-1}_{t}}=\alpha_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}$ and $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$, we have
$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
This concludes the proof.$\square$
D.4 Proof of Proposition 5
This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3. Without loss of generality, we let $M>\lambda_{2}$. Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$$\square$
D.5 Proof of Theorem 3
Define $x^{\dagger}(\theta)\coloneqq\theta/\|\theta\|$. The conclusion of Theorem 2 holds for $x^{\dagger}(\theta)$, i.e. for all $t\in[1,T]$ and all $s\leq t$, we have $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s}$ and $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}$.
For any $\theta$, the feasible optimal action $x^{\star}_{s}(\theta)$ is the point in $\mathcal{X}_{s}$ that has the smallest distance from $x^{\dagger}(\theta)$. Then we have $\|x^{\dagger}(\theta)-x^{\star}_{s}(\theta)\|\leq\rho_{s}$ by the definition of the Euclidean covering radius. Therefore, let $\theta_{0}$ denote the $\theta$ that attains $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}$, it holds that
$\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}=\|x^{\star}_{s}(\theta_{0})\|_{V_{s}^{-1}}$ $\geq\|x^{\dagger}(\theta_{0})\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}).$
Similarly, we have $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$. As a result, it holds that $\alpha_{s}\leq(\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))/(\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))$$\square$
Appendix E TS-MR and Greedy-MR Algorithms
Algorithm 2 TS-MR (Greedy-MR)
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
     $\widehat{\mu_{t}}\leftarrow\widehat{\alpha}_{t}(1+\iota_{t})+1+\iota_{t}$
     if $\widehat{\mu}_{t}\leq\mu$ then
        Sample $\eta_{t}\sim\mathcal{D}^{SA}(\delta^{\prime})$ (defined in Section 3)
        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     else
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     end if
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
  end for

### v3

@@
-Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits
-Abstract
-This paper is motivated by recent developments in the linear bandit literature, which have revealed a discrepancy between the promising empirical performance of algorithms such as Thompson sampling and Greedy, when compared to their pessimistic theoretical regret bounds.
-The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances.
-To address this, we propose a new data-driven technique that tracks the geometry of the uncertainty ellipsoid, enabling us to establish an instance-dependent frequentist regret bound for a broad class of algorithms, including Greedy, OFUL, and Thompson sampling.
-This result empowers us to identify and “course-correct” instances in which the base algorithms perform poorly.
-The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$, while retaining most of the desirable properties of the base algorithms.
-We present simulation results to validate our findings and compare the performance of our algorithms with the baselines.
+
+[…]
+[…]
 1 Introduction
-Multi-armed bandits (MABs) have garnered significant attention as they provide a well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems.
+Multi-armed bandits (MABs) have garnered significant attention as they provide well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems.
 In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards.
-This paper specifically focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases.
-In this model, the system parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield rewards equal to the inner product of $\theta^{\star}$ and the chosen action, along with mean-zero, subgaussian noise.
+This paper focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases.
+In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with mean equal to the inner product of $\theta^{\star}$ and the chosen action.
 The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time.
-The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total rewards achieved by the policy and the maximum achievable reward.
+The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total expected rewards achieved by the policy and the maximum achievable expected reward.
 Achieving this objective necessitates striking a balance between exploration and exploitation.
 In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards.
-Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985).
-Other popular strategies include Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions.
+Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010).
+Another popular strategy is Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions.
 Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
@@
 Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al.
-(2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to logarithmic factors.
-On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of noise and algorithm.
-In the frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al.
+(2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to polylogarithmic factors.
+On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of the noise and the algorithm.
+In this frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al.
 (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound.
-However, for a frequentist variant of LinTS referred to as TS-Freq, Agrawal and Goyal (2013) and Abeille et al.
-(2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret by inflating the posterior distribution.
+However, for a frequentist variant of LinTS referred to as TS-Freq, which inflates the posterior distribution, Agrawal and Goyal (2013) and Abeille et al.
+(2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret.
 This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$.
-Recent work by Hamidi and Bayati (2020a) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved.
-While no general theoretical guarantee exists for the Greedy algorithm in linear bandit problems, it is known that both Greedy and LinTS perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
-Despite the theoretical gaps in performance, LinTS has shown strong empirical performance, indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently.
+Recent work by Hamidi and Bayati (2020b) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved.
+For the Greedy algorithm in linear bandit problems, no general theoretical guarantee exists.
+Moreover, it is known that both LinTS and Greedy might perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
+Despite the theoretical gaps in performance, LinTS has shown strong empirical performance (Russo et al., 2018), indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently.
 Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021).
@@
 (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency.
-This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify problematic instances where LinTS and Greedy could potentially fail, and apply a “course-correction” to ensure competitive frequentist regret bounds?
+This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify, in a data-driven fashion, problematic instances where LinTS and Greedy could potentially fail and apply a “course-correction” to ensure competitive frequentist regret bounds?
 And can this be achieved without compromising their impressive empirical performance and computational efficiency?
@@
 Specifically, we make the following contributions.
-  • Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid.
-This technique is essential in maximizing the utilization of historical data, thereby enabling the “course-correction”.
-  • We derive a general instance-dependent frequentist regret bound for the proposed POFUL algorithms, a family of algorithms including OFUL, LinTS, TS-Freq, and Greedy as special cases.
-The bound can be efficiently calculated in a data-driven manner based on the observation history.
-  • We propose variants of LinTS and Greedy that enjoy the minimax optimal frequentist regret, while retaining most of the desirable properties such as computational efficiency.
+  • We develop a real-time geometric analysis technique for the $d$-dimensional confidence ellipsoid surrounding $\theta^{\star}$.
+This method is crucial for maximizing the use of historical data, advancing beyond methods that capture only limited information from the confidence ellipsoid, such as a single numerical value.
+Consequently, this facilitates a more precise “course-correction”.
+  • We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, instance-dependent frequentist regret bound for them.
+This bound is efficiently computable using data observed from previous decision epochs.
+  • We introduce course-corrected variants of LinTS and Greedy that achieve minimax optimal frequentist regret.
+These adaptations maintain most of the desirable characteristics of the original algorithms.
 The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid.
-This is in stark contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their overly narrow concentration on comparatively rudimentary summary statistics.
-Related literature.
-The ideas presented in this paper bear similarities to existing literature that explores different forms of exploration (Russo and Van Roy, 2016; Kirschner and Krause, 2018) or recent works by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b) that aim to utilize observed data for reducing exploration.
-However, the methodology and the manner in which the data is utilized in this paper differ significantly.
-For instance, Bastani et al.
-(2021) examines the minimum eigenvalue of the sample covariance matrix which is a single-parameter summary statistic of the observed data, while Hamidi and Bayati (2020b) incorporates information from the (one-dimensional) confidence intervals around the reward values.
-In contrast, our approach leverages the complete geometric information of the $d$-dimensional confidence ellipsoid.
+This is in contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their focus on summary statistics.
+1.1 Other Related literature
+We note that Valko et al.
+(2014); Kocák et al.
+(2014, 2020); Kocák and Garivier (2020) consider the spectral of graphs and study Spectral Bandits.
+In their work, the rewards of arms are represented as smooth functions on a graph, allowing the exploitation of a low-rank structure for improved algorithmic performance and a regret guarantee that is independent of the original number of actions.
+In contrast, our approach leverages the spectral properties of the sample covariance of actions, which is distinct from graph spectral analysis.
+Additionally, our research addresses the broader context of stochastic linear bandit problems without presupposing any low-rank structure.
+The concepts in this paper bear similarities to existing research that investigates various exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018), as well as recent studies by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a) that focus on using observed data for reducing exploration.
+Nonetheless, our methodology and the specific use of data markedly differ.
+For example, Bastani et al.
+(2021) emphasizes the minimum eigenvalue of the sample covariance matrix, essentially a single-parameter summary of the observed data.
+On the other hand, Hamidi and Bayati (2020a) uses information from the one-dimensional confidence intervals surrounding the reward values.
+A more closely related paper, Hamidi and Bayati (2020b), uses spectral information to enhance the performance of TS in LB problems.
+They look at a single summary statistics of the confidence set that they call thinness coefficient, to decide if the posterior of TS should be inflated.
+In contrast, with these studies, our approach capitalizes on the full geometric details of the $d$-dimensional confidence ellipsoid, utilizing richer geometric information.
+1.2 Organization
+The structure of this paper is as follows: Section 2 introduces the notations and problem formulation.
+In Section 3-4, we present the POFUL family of algorithms and analyze their frequentist regret.
+Sections 5-6 detail our data-driven approach and meta-algorithm for course-correction.
+Section 7 discusses simulations using synthetic and real-world datasets.
+The paper concludes in Section 8 with a summary of findings.
+All technical proofs, further discussions, and supplementary materials are provided in the appendices.
+endcomment
 2 Setup and Preliminaries
@@
 The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$.
-Without loss of generality, we normalize the actions such that $\mathcal{X}_{t}\subset{\mathcal{S}}_{d-1}$ for all $t\in[T]$.
 Assumption 3 (Noise)
@@
 In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator.
-Fix a regularization parameter $\lambda$, the design matrix, and the Regularized Least Square (RLS) estimate are respectively given by
+Fix a regularization parameter $\lambda$, the sample covariance matrix, and the Regularized Least Square (RLS) estimate are respectively given by
 $V_{t}=\lambda I+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
@@
 Proposition 1 (Theorem 2 in Abbasi-Yadkori et al. (2011))
-Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that
-$\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta),\quad\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$
-for all $x\in\mathbb{R}^{d}$, where
+Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that $\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)$ and $\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$ for all $x\in\mathbb{R}^{d}$, where
 $\beta_{t}^{RLS}(\delta)=R\sqrt{2\log\frac{(\lambda+t)^{d/2}\lambda^{-d/2}}{\delta}}+\sqrt{\lambda}S.$ (2)
@@
 The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret.
-Basically, this proposition is saying that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
+Essentially, this proposition suggests that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
 Proposition 2 (Lemma 11 in Abbasi-Yadkori et al. (2011))
@@
 3 POFUL Algorithms
-In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL).
+In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). It allows us to analyze the frequentist regret of common algorithms in a unified way.
+At a high level, POFUL is designed to encompass the exploration mechanism of OFUL and LinTS.
 POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$.
@@
 A pseudocode representation of POFUL is provided in Algorithm 1.
+Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid.
+We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$.
+By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
 Algorithm 1 POFUL
-0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
+0:  $T$, $\delta$, $\lambda$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
   Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
-  for t = 1, 2 , …, T  do
+  for t = 0, 1, …, T  do
      Sample a feasible pivot $\widetilde{\theta}_{t}$ with respect to $\iota_{t}$ according to Definition 2
@@
   end for
-The intuition behind POFUL is combining the exploration mechanism of OFUL and LinTS.
-OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores by sampling from the confidence ellipsoid.
-We allow POFUL to select an arbitrary pivot from $\mathcal{E}_{t}^{PVT}$ to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$.
-By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
 OFUL
@@
 On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al.
-(2017), denoted by TS-Freq.
+(2017), namely TS-Freq.
 This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
@@
 For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$.
-However, later we see this could be problematic when bounding $\mathcal{R}^{PE}(T)$.
+However, later we see this could be problematic for bounding $\mathcal{R}^{PE}(T)$.
 $\mathcal{R}^{PE}(T)$ corresponds to the pessimism term.
@@
 (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e.
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$.
+$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ with a constant probability.
 For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures.
-The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy.
+The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy that cannot be analyzed by the optimism-based methods.
 In fact, non-optimistic samples might be able to control the pessimism term.
@@
 In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$.
-The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix D.2.
+The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix 10.2.
 Proposition 4
@@
 $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
-On the RHS, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret.
-To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$ norm.
+On the right-hand side, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret.
+To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$-norm.
+This motivates the following definition.
 Definition 5
 For each $t\geq 1$, let $\widetilde{x}_{t}$ and $x_{t}^{\star}$ respectively denote the action chosen by POFUL and the optimal action.
-We define the sensitivity ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$.
-We also define the regret ratio at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
-Recall the Cauchy–Schwarz inequality $\langle{x},{\theta_{1}-\theta_{2}}\rangle\leq\|x\|_{V_{t}^{-1}}\|\theta_{1}-\theta_{2}\|_{V_{t}}$ , intuitively $\alpha_{t}$ serves as a measure of how sensitive the reward $\langle{\widetilde{x}_{t}},{\theta}\rangle$ is compared with that of the optimal action $\langle{x^{\star}_{t}},{\theta}\rangle$, with respect to the change of the parameter $\theta$ in the $V_{t}$-norm.
-The intuition behind the definition for $\mu_{t}$ is connecting POFUL and OFUL.
-It provides an upper bound for the ratio of instantaneous regrets incurred by POFUL and the OFUL-type upper bound.
+We define the uncertainty ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$.
+We also define the (instantaneous) regret proxy at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
+Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$.
+Hence, $\alpha_{t}$ serves as the ratio of uncertainty degrees of the reward obtained by the optimal action $x^{\star}_{t}$ and the chosen action $\widetilde{x}_{t}$.
+The intuition behind the definition for $\mu_{t}$ is constructing a regret upper bound similar to that of OFUL.
 Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$.
-In this sense, our definition of $\mu_{t}$ and $\alpha_{t}$ can be extended to OFUL.
-One feasible choice is setting $\mu_{t}\equiv 2$ for OFUL and therefore $\alpha_{t}=(\mu_{t}-1-\iota_{t}-\tau_{t})/(1+\iota_{t}-\tau_{t})=1$ (recall $\iota_{t}=0$ and $\tau_{t}=1$ for OFUL).
-The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix D.3.
+In this sense, $\mu_{t}$ is a proxy of the instantaneous regret incurred by POFUL at time $t$.
+Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}\equiv 1$ for OFUL (recall $\iota_{t}\equiv 0$ and $\tau_{t}\equiv 1$ for OFUL).
+The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix 10.3.
+It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
 Theorem 1 (Oracle frequentist regret bound for POFUL)
@@
 Remark 1
-A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al.
-(2017), where they indicate their proof of regret bound can be reproduced if the algorithm samples $\widetilde{\theta}_{t}$ that satisfies $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$.
-They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation.
-Our work answers this question positively in the next section by investigating the property of (the reciprocal of) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
-Remark 2
 We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$.
@@
 Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy.
-Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL.
-This is indeed our aim in the next section.
-5 A Data-Driven Approach to Calibrating POFUL
+Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL, and could be utilized later for the course correction.
+This is indeed our aim in the next section and we achieve it using geometric arguments.
+Remark 2
+A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al.
+(2017), where they indicate their proof of regret bound for LinTS can be reproduced if LinTS samples $\widetilde{\theta}_{t}$ satisfying $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$.
+They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation.
+Our work answers this question positively in the next section by investigating the property of (the reciprocal of their) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
+5 A Data-Driven Approach
 In this section, we present the main contribution of this work which provides a data-driven approach to calibrating POFUL.
-We’ll focus on LinTS and Greedy as they lack theoretical results of frequentist regret bound.
-In these settings, we have $\tau_{t}=0$ and hence $\mu_{t}\coloneqq(1+\alpha_{t})(1+\iota_{t})$.
-Note that $\iota_{t}$ is the inflation parameter that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound for the sensitivity ratio $\alpha_{t}$, which turns into an upper bound for the regret ratio $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
-In the following, we first construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$.
-Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids.
-Subsequently, we show how these bounds can be utilized to calibrate POFUL.
-Specifically, we propose the TS-MR and Greedy-MR algorithms that attain a provable minimax-optimal frequentist regret up to constant factors, and yet keep the majority of their original properties such as computational efficiency and low regret in most instances.
-5.1 Upper Bounding the Sensitivity Ratio
-We consider two scenarios: (1) Continuous action space and (2) Discrete action space.
-For sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for the first scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al.
+Note that $\iota_{t}$ and $\tau_{t}$ are parameters of POFUL that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$, which turns into an upper bound $\widehat{\mu}_{t}$ for the regret proxy $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
+Typically, we can check $\widehat{\alpha}_{t}$ is no less than one because we might have $\widetilde{x}_{t}=x_{t}^{\star}$.
+As a result, it holds that $\widehat{\mu}_{t}=(1+\widehat{\alpha}_{t})(1+\iota_{t})+(1-\widehat{\alpha}_{t})\tau_{t}\geq(1+\widehat{\alpha}_{t})(1+\iota_{t})$, i.e., non-zero $\tau_{t}$ gives a larger regret proxy.
+Given this observation, we’ll focus on scenarios where $\tau_{t}\equiv 0$.
+This avoids the large regret proxies and computational inefficiencies seen in OFUL when determining the action $\widetilde{x}_{t}$.
+Such scenarios include LinTS and its variants like TS-Freq as well as Greedy - all of which are standard algorithms that still lack theoretical regret guarantee results.
+In the following, we construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for both discrete and continuous action scenarios.
+5.1 Discrete Action Space
+We start with the discrete action space scenario to illustrate our method.
+It turns out that upper bounding $\alpha_{t}$ is done by simply comparing the confidence interval lengths of all potentially optimal actions.
+To see this, recall for an arbitrary action $x\in\mathcal{X}_{t}$, by Proposition 1, with high probability its expected reward is bounded as
+$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\eqqcolon U_{t}(x).$
+The confidence interval length is proportional to $\|x\|_{V_{t}^{-1}}$.
+Therefore, to characterize the range of $\|x_{t}^{\star}\|_{V_{t}^{-1}}$, we only need to identify the set of potentially optimal actions and calculate the minimal $V_{t}^{-1}$-norm among all actions in that set.
+Note that for an action $x\in\mathcal{X}_{t}$ to remain potentially optimal, it cannot be dominated by another action.
+I.e., we require $U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}L_{t}(y)$.
+Therefore, define
+$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\},$
+the set of potentially optimal actions is given by $C_{t}(\beta_{t}^{RLS}(\delta^{\prime}))$.
+Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))$, where we let $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$.
+The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
+By comparing the range of $V_{t}^{-1}$-norm for both sets, an upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$ is given by
+$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (5)
+5.2 Continuous action space.
+When it comes to the generic settings where the number of actions is large or when the set of actions is a continuous set in $\mathbb{R}^{d}$, the upper bound above is inefficient or even infeasible to compute.
+In this section, we tackle this issue.
+Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids, providing upper bounds that can be computed efficiently.
+For the sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for this scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al.
 (2017).
-The analysis of the second scenario is deferred to Appendix A.
-Continuous action space.
-Note that in this case, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$, we only need to study the property of the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$.
-A direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the ellipsoids.
+We remark that for this specific setting, the problem is still hard.
+This is because we don’t have a closed-form solution for the set of potentially optimal actions.
+In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$.
+To upper bound $\alpha_{t}$, a direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$.
 Specifically, we have
 $\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (6)
-At first glance, one might think that (5) is too conservative to provide useful calibration for POFUL.
-However, it turns out that the lower bound is pretty tight in practice.
 As is illustrated in Figure 1.
-The intuition is that, when POFUL has implemented sufficient exploration so that the confidence ellipsoid $\mathcal{E}_{t}$ is small enough, the potentially optimal actions concentrate in a small cap $\mathcal{C}_{t}$ on the unit hypersphere ${\mathcal{S}}_{d-1}$.
-All actions within $\mathcal{C}_{t}$ roughly point to the same direction and thus have similar $V_{t}^{-1}$-norm.
-The following main theorem of this paper derives a upper bound of $\alpha_{t}$ based on this idea.
-We sketch the proof in Appendix B.
+The set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
+It’s hard to get a closed-form expression for $\mathcal{C}_{t}$, so we cannot directly calculate the range of $V_{t}^{-1}$-norm of actions in $\mathcal{C}_{t}$.
+Nevertheless, we can approximate the range by investigating the geometric properties of the ellipsoids.
+The intuition here is that, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$.
+All actions within $\mathcal{C}_{t}$ point to similar directions and thus have similar $V_{t}$-norm.
+(Note that for a unit vector $x$, we have that $\|x\|_{V_{t}}$ can be written as a weighted summation of the eigenvalues of $V_{t}$, where the weights are determined by the direction of $x$.) Hence we can bound the range of $V_{t}$-norm by investigating the concentration of $\mathcal{C}_{t}$, which is pure geometry, and we’ll use the range of $V_{t}$-norm to determine the range of $V_{t}^{-1}$-norm.
 Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$.
-The following assumption simplifies the proof, and we remark that it holds almost surely.
-Assumption 4
-For all $t$, $V_{t}$ has distinct eigenvalues, i.e. $\lambda_{1}(V_{t})>\lambda_{2}(V_{t})\dots>\lambda_{d}(V_{t})\geq 0$.
+(a): $\mathcal{C}_{t}$ is the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
+(b): As the decision-maker collects more and more data, the confidence ellipsoid $\mathcal{E}_{t}$ shrinks accordingly (here different color corresponds to different levels of exploration).
+The potentially optimal actions point to similar directions, which essentially determine their $V_{t}$-norm.
+Hence it’s plausible one could get a proper range of their $V_{t}$-norm using geometry.
+The following main theorem of this paper derives an upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Section 5.3.
 Theorem 2
-Under Assumpsion 4, define $\beta^{PVT}_{t}=\iota_{t}\beta^{RLS}_{t}$ and
+Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$. Define $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$ and
 […]
@@
 Case 2: linear structures.
-When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$.
+When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated in Appendix 11).
 To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small.
-The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding sensitivity ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1.
+The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1.
 This provides a plausible explanation for the empirical success of LinTS and Greedy.
@@
 Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{\max}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
-5.2 TS-MR and Greedy-MR Algorithms
-In the following, we propose the algorithm Linear Thompson Sampling with Maximum (Regret) Ratio (TS-MR).
-The intuition is to detect the performance of LinTS by estimating $\mu_{t}$ and avoid bad LinTS actions by switching to OFUL occasionally.
-Specifically, at each time $t$, TS-MR calculates an estimate $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$.
+Remark 3
+Clearly, one can further improve the bound in Equation (6) by plugging in the selected action $\widetilde{x}_{t}$.
+This is true for standard algorithms such as OFUL, LinTS, and Greedy.
+Explicitly, the uncertainty ratio can be bounded using $\widetilde{x}_{t}$ and Theorem 2 as
+$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}\leq\frac{\Phi_{t}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}.$ (7)
+However, for the variants of POFUL that we are going to propose in the Section 6, it is necessary to calculate the bound before selecting the action $\widetilde{x}_{t}$.
+Therefore, the bound shouldn’t depend on $\widetilde{x}_{t}$.
+5.3 Sketch of the Proof
+In this section, we present the proof of Theorem 2.
+In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
+First note that, when $\|\widehat{\theta}_{t}\|_{V_{t}}<\beta_{t}$, the bound in Theorem 2 becomes
+$\alpha_{t}\leq\lambda_{d}^{-\frac{1}{2}}(V_{t})/\lambda_{1}^{-\frac{1}{2}}(V_{t})=\sqrt{\lambda_{1}(V_{t})/\lambda_{d}(V_{t})}.$
+This bound holds trivially using the fact that $\lambda_{1}(V_{t})\leq\|x\|^{2}_{V_{t}}\leq\lambda_{d}(V_{t})$.
+This is the case when the data is insufficient and the confidence interval is too large to get a non-trivial upper bound for $\alpha_{t}$.
+In the following, without loss of generality we assume $\|\widehat{\theta}_{t}\|_{V_{t}}^{2}\geq\beta_{t}^{2}$.
+The proof decomposes into three steps.
+In the first two steps, as is illustrated in $\mathbb{R}^{2}$ in Figure 2, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below.
+Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid.
+Next, we show that upper and lower bounds for the $V_{t}$-norm can be converted into upper and lower bounds for the $V_{t}^{-1}$-norm by solving a linear programming problem.
+Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound.
+We sketch the proof below and postpone the detailed proof for all lemmas in this section to Appendix 9.
+Figure 2: Illustration of Step 1 and 2 in $\mathbb{R}^{2}$.
+Orange dashed rays: rays starting from the origin might have different numbers of intersections with $\mathcal{E}_{t}$, indicating whether the corresponding action lies in the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
+Blue dashed curve: the ellipsoid with fixed $V_{t}$-norm $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$.
+The intersection of this ellipsoid and $\mathcal{E}_{t}$ has the same projection as $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
+In the following, we let $\phi_{t}\coloneqq\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-\beta_{t}^{2}}$, which is well-defined since we assume $\|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta_{t}$.
+In geometry, one can show that for a ray starting from the origin and intersecting $\mathcal{E}_{t}$ only at one point, $\phi_{t}$ is the ${V_{t}}$-norm of the intersection point.
+Step 1. Upper Bounding the $V_{t}$-norm of actions.
+Our first lemma investigates such intersection and provides an upper bound for the $V_{t}$-norm for the optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$.
+The proof is based on investigating the condition for a ray paralleling an action $x\in{\mathcal{S}}_{d-1}$ to intersect with $\mathcal{E}_{t}$, which means $x$ is in the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$ and might become the optimal action $x^{\star}(\theta)$.
+Lemma 1
+For any $\theta\in\mathcal{E}_{t}$, we have $\|x^{\star}(\theta)\|_{V_{t}}\leq\|\widehat{\theta}_{t}\|_{V_{t}^{2}}/\phi_{t}$.
+Step 2. Lower Bounding the $V_{t}$-norm of actions.
+In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$.
+$\mathcal{H}_{t}$ consists of $\theta\in\mathcal{E}_{t}$ whose $V_{t}$-norm is $\phi_{t}$.
+One can check $\mathcal{H}_{t}$ is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$, and the projection of $\mathcal{H}_{t}$ onto ${\mathcal{S}}_{d-1}$ is the same as that of $\mathcal{E}_{t}$ by convexity.
+Hence, it suffices to only consider $\theta\in\mathcal{H}_{t}$ as the corresponding set of optimal actions coincides.
+A lower bound for the $V_{t}$-norm is given by the following lemma.
+Lemma 2
+For any $\theta\in\mathcal{E}_{t}$, we have
+$\|x^{\star}(\theta)\|_{V_{t}}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
+The proof is directly using the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$.
+Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
+Step 3. Bounding the $V_{t}^{-1}$-norm of actions
+The following lemma determines the range of action $x$’s $V_{t}^{-1}$-norm based on its $V_{t}$-norm range.
+It turns out that the two ranges can be related using the spectral information of the sample covariance matrix $V_{t}$, which is related to the shape of the confidence ellipsoid.
+Lemma 3
+Let $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{V}}\}$ be the set of distinct eigenvalues of $V$ such that $\lambda_{1}>\lambda_{2}>\dots>\lambda_{N_{V}}>0$.
+Let $x\in\mathcal{B}_{d}$ satisfies $0<m\leq\|x\|_{V}^{2}\leq M$.
+We have
+$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}},$ (8)
+where $k$ is such that $\lambda_{k}=\max_{i\in[N_{v}]}\{\lambda_{i}\geq M\}$.
+The proof involves expressing the $V$- and $V^{-1}$-norms as weighted sums of $V$’s eigenvalues, then solving a linear programming (LP) problem constrained by the norm ranges.
+By inserting the upper and lower bounds of the $V_{t}$-norm from Lemmas 1 and 2 into Lemma 3, we finalize the proof of Theorem 2.
+6 A Meta-Algorithm for Course-Correction
+In this section, we show how the data-driven regret bound can be utilized to improve standard bandit algorithms.
+Specifically, we propose a meta-algorithm that constructs course-corrected variants of the base algorithms.
+This ensures they achieve a minimax-optimal frequentist regret guarantee (up to constant factors) while retaining the majority of the original characteristics of the base algorithms, such as computational efficiency and low regret in most cases.
+We take LinTS as an example and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR).
+The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions when necessary.
+Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$.
 If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action.
 We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm.
-The pseudocode is presented in Appendix E.
-By construction, TS-MR and Greedy-MR have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
-Corollary 1
-TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
-6 Simulations
-(a) example 1
-(b) example 2
-(c) example 3
-Figure 2: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
-In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
-Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
-We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm.
-At each time $t$, we generate 100 i.i.d.
-random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$.
-This is a basic example of standard stochastic linear bandit problems without any extra structure.
-We set the threshold $\mu=8$ for TS-MR and Greedy-MR.
-We average simulation results over 20 independent runs.
-The results are summarized in Figure 1(a).
-TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
-Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
-We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$.
-At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0.
-Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks.
-This problem is equivalent to a 10-armed contextual bandit.
-We ser $\mu=12$ for TS-MR and Greedy-MR.
-Again, we average simulation results over 20 independent runs.
-The results are summarized in Figure 1(b).
-In this setting, Greedy performs suboptimally due to a lack of exploration for all arms.
-Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects bad regret ratio $\widehat{\mu}_{t}$.
-Example 3. Prior mean mismatch (Hamidi and Bayati, 2020a).
-This is an example in which LinTS is shown to incur linear Bayesian regret.
-We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$.
-It is shown in Hamidi and Bayati (2020a) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$.
-Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action.
-We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit.
-We ser $\mu=12$ for TS-MR and Greedy-MR.
-We carried out 20 independent runs.
-The results are summarized in Figure 1(c).
-We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
-7 Conclusion
-In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, and Greedy as special cases.
-Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm.
-Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest.
-As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail.
-We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
-Appendix A Discrete Action Space
-In this scenario, we consider upper bounding the sensitivity ratio for discrete action space.
-Compared with the unit hypersphere ${\mathcal{S}}_{d-1}$ considered in Section 5.1 , the main difference is that the oracle optimal action $\theta/\|\theta\|$ for $\theta$ might not lie in the action space $\mathcal{X}_{t}$, so we don’t have a closed form for the attainable optimal action $x^{\star}_{t}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$.
-To tackle this problem, we need to characterize the property of all potentially optimal actions.
-We provide two approaches to solve this problem.
-This first approach works for the setting where the cardinalities of the action sets are relatively small.
-The second approach utilizes the statistical properties of the action sets and works for the setting where the cardinalities of the action sets are large.
-Small cardinality setting.
-Note that the optimistic reward of POFUL when $\tau_{t}=0$ is $\langle{x},{\widetilde{\theta}_{t}}\rangle$, where $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$ is an arbitrary pivot parameter.
-We have $L_{t}(x)\leq\langle{x},{\widetilde{\theta}_{t}}\rangle\leq U_{t}(x)$, where we define the confidence lower and upper bounds as
-$L_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}),\quad U_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$
-The potentially optimal $x$ that might optimize $\langle{x},{\widetilde{\theta}_{t}}\rangle$ should satisfy
-$U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}\ L_{t}(y),$
-i.e. $x$ is not dominated by any other action in $\mathcal{X}_{t}$. Therefore, define
-$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\}.$
-An upper bound for $\alpha_{t}$ can be calculated as
-$\alpha_{t}\leq\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (6)
-Large cardinality setting.
-In this setting, we characterize the distance between $\theta/\|\theta\|$ and $x^{\star}_{t}(\theta)$. We first introduce the notion of Euclidean covering radius.
-Definition 6
-The Euclidean covering radius $\rho_{t}$ of a set $\mathcal{X}_{t}\subseteq{\mathcal{S}}_{d-1}$ is given by
-$\rho_{t}=\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\coloneqq\max_{y\in{\mathcal{S}}_{d-1}}\min_{x\in\mathcal{X}_{t}}\|y-x\|.$
-With the help of the Euclidean covering radius, we prove the following variant of Theorem 2. The proof is postponed to Appendix D
-Theorem 3
-Under Assumpsion 4, let $\{\rho_{t}\}$ be the sequence of Euclidean covering radius of $\mathcal{X}_{t}$, define $\Psi_{t},\Phi_{t}$ as in Theorem 2.
-Then for all $t\in[1,T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds that
-$\alpha_{s}\leq\frac{\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}{\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}$ (7)
-for all $s\leq t$.
-To justify this new lower bound, it remains to show that $\rho_{t}$ is small enough under proper conditions.
-For the sake of a better illustration, we focus on action sets that consist of uniformly and independently distributed points on ${\mathcal{S}}_{d-1}$.
-The following proposition shows that the Euclidean covering radius has an high-probability upper bound that decays at the order of $\widetilde{\mathcal{O}}(N^{-1/d+o(1)})$, where $N$ is the cardinality of actions.
-Proposition 6 (Corollary 2.9 in Reznikov and Saff [2016])
-Suppose $\mathcal{X}$ is a set of $N$ uniformly and independently distributed random points on the unit sphere ${\mathcal{S}}_{d-1}$, for any $\kappa>0$, there exist constants $C_{\kappa,d}>0$, such that
-$\mathbb{P}\left[\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\leq C_{\kappa,d}\left(\frac{\log N}{N}\right)^{1/d}\right]\geq 1-N^{-\kappa}.$
-Appendix B Sketch of the Proof of Theorem 2
-In this section, we present the proof of Theorem 2.
-In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
-The proof of Theorem 2 decomposes into two steps.
-First, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below.
-Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid.
-Second, we show that upper and lower bounds for the $V_{t}$-norm can be transferred into upper and lower bounds for the $V_{t}^{-1}$-norm.
-Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound.
-We postpone the detailed proof for all lemmas in this section to Appendix C.
-Step 1. Bounding the $V_{t}$-norm of actions.
-We illustrate our geometric approach in $\mathbb{R}^{2}$ in Figure 3.
-We first upper bound the $V_{t}$ norm for optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$.
-In the following, we define $\phi_{t}=\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}-\beta_{t}^{2}}$.
-Lemma 1
-For any $\theta\in\mathcal{E}_{t}$, we have $\|x^{\star}(\theta)\|_{V_{t}}\leq\|\widehat{\theta}_{t}\|_{V_{t}^{2}}/\phi_{t}$.
-The proof is based on the observation that for any $\theta\in\mathcal{E}_{t}$, the line $\{t\theta:t\in\mathbb{R}\}$ (orange dashed lines in Figure 3) and ellipsoid $\mathcal{E}_{t}$ (in green) must have at least one intersection point.
-In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$ (blue dashed curve).
-We can check this intersection is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$.
-A lower bound for the $V_{t}$-norm is given by the following lemma.
-Lemma 2
-For any $\theta\in\mathcal{H}_{t}$, we have
-$\|x^{\star}(\theta)\|_{V_{t}}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
-The proof is based on the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$.
-Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
-Figure 3: Illustration of Step 1 in $\mathbb{R}^{2}$.
-Step 2. Bounding the $V_{t}^{-1}$-norm of actions
-The following lemma bounds the $V_{t}^{-1}$-norm of an action $x$ given a range of its $V_{t}$-norm.
-Lemma 3
-Suppose $V$ has distinct eigenvalues $\lambda_{1}(V)>\lambda_{2}(V)\dots>\lambda_{d}(V)$ and $x\in\mathcal{B}_{d}$ satisfies $m\leq\|x\|_{V}^{2}\leq M$, where $M\geq 0$ is such that $\lambda_{k}(V)\leq M<\lambda_{k+1}(V)$.
-We have
-$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$ (8)
-The proof is based on writing $x$ as a linear combination of the eigenvectors of $V_{t}$, $x=\sum_{i=1}^{d}w_{i}u_{i}$, and convert the original problem into a linear programming problem with constraints using the expressions $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
-Combining the two steps, we complete the proof of Theorem 2.
-Appendix C Proof of Lemmas for Theorem 2
-C.1 Proof of Lemma 1
-Let $\theta=tx$ where $x$ is any unit vector in $\mathbb{R}^{d}$ and $t\in\mathbb{R}$ is a scalar.
-Consider the equation that characterizes the intersection $\{tx:t\in\mathbb{R}\}\cap\mathcal{E}_{t}$, namely $(tx-\widehat{\theta}_{t})^{\top}V_{t}(tx-\widehat{\theta}_{t})=\beta_{t}^{2}$.
-Equivalently, we have $t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=0$.
-This quadratic equation of $t$ has at least one root if the discriminant is non-negative, i.e.
-$4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$.
-Then by direct computation,
-$\|x\|_{V_{t}}\leq\frac{\sqrt{(x^{\top}V_{t}\widehat{\theta}_{t})^{2}}}{\phi_{t}}\leq\frac{\sqrt{x^{\top}x}\sqrt{(\widehat{\theta}_{t})^{\top}V_{t}^{\top}V_{t}\widehat{\theta}_{t}}}{\phi_{t}}=\frac{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}}{\phi_{t}}.$
-Note that $\theta\in\mathcal{E}_{t}$ if and only if $\theta=tx$ for some $x$ that satisfies $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, we conclude the proof. $\square$
-C.2 Proof of Lemma 2
-Note that for any $\theta\in\mathcal{H}_{t}\subset\mathcal{E}_{t}$, it holds that $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$.
-Also, by the construction of $\mathcal{H}_{t}$, we have $\|\theta\|_{V_{t}}=\phi_{t}$.
-Then we conclude the proof by the definition
-$\|x^{\star}(\theta)\|_{V_{t}}=\frac{\|\theta\|_{V_{t}}}{\|\theta\|}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
-$\square$
-C.3 Proof of Lemma 3
-Let $V=U\Lambda U^{\top}$ be the eigendecomposition of $V$, so that the columns of $U$, $\{u_{1},\dots,u_{d}\}$, form an orthonormal basis of $\mathbb{R}^{d}$ and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$ is a diagonal matrix.
-We can write $x$ as a linear combination of $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$: $x=\sum_{i=1}^{d}w_{i}u_{i}$, where $\sum_{i=1}^{d}w_{i}^{2}=1$.
-By direct computation, we have $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
-We would like to study the range of $\|x\|_{V^{-1}}^{2}$ when $\|x\|_{V}^{2}$ is bounded as $m\leq\|x\|_{V}^{2}\leq M$.
-First, let’s focus on maximizing $\|x\|_{V^{-1}}^{2}$, it suffices to solve the LP problem
-maximize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}\quad\text{s.t.}\quad\forall i,a_{i}\geq 0,\ \sum_{i=1}^{d}a_{i}=1,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\leq M.$
-The Lagrangian is given by
-$L(d,\mu,\eta)=\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}+\mu(1-\sum_{i=1}^{d}a_{i})+\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)+\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})+\sum_{i=1}^{d}\kappa_{i}a_{i}.$
-The KKT conditions are given by
-$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{d}a_{i}=1,\sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{d}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$
-Since $\lambda_{i}$’s are distinct, to satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices.
-Hence, $a_{i}$ can only be non-zero for at most two indices, denoted by $a_{i}$ and $a_{j}$.
-We write $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$ and hence $\lambda_{i}>\lambda_{j}$.
-Let $y=\|x\|_{V}^{2}\in[\lambda_{j},\lambda_{i}]$, we have $a\lambda_{i}+(1-a)\lambda_{j}=y$ and $a\lambda_{i}^{-1}+(1-a)\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
-By direct computation, $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$, and clearly,
-$\begin{cases}\frac{\partial f}{\partial y}&=-\frac{1}{\lambda_{i}\lambda_{j}}<0,\\ \frac{\partial f}{\partial\lambda_{i}}&=(\frac{y}{\lambda_{j}}-1)\frac{1}{\lambda_{i}^{2}}>0,\\ \frac{\partial f}{\partial\lambda_{j}}&=(\frac{y}{\lambda_{i}}-1)\frac{1}{\lambda_{j}^{2}}<0.\end{cases}$
-Then the maximum of $\|x\|_{V^{-1}}^{2}$ is obtained by setting $\lambda_{i}=\lambda_{1}$, $\lambda_{j}=\lambda_{d}$, $y=m$, and we have
-$x^{*}_{\max}$ $=\frac{m-\lambda_{d}}{\lambda_{1}-\lambda_{d}}\lambda_{1}+\frac{\lambda_{1}-m}{\lambda_{1}-\lambda_{d}}\lambda_{d},\quad\|x^{*}_{\max}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$
-Similarly, by investigating the KKT conditions for the LP problem that minimize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}$, the minimum of $\|x\|_{V^{-1}}^{2}$ is obtained by some $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$, $\lambda_{i}>\lambda_{j}$.
-Setting $\lambda_{u}=\min\{\lambda_{i}\geq M\}$, $\lambda_{l}=\max\{\lambda_{i}\leq M\}$, $y=M$, and we have
-$x^{*}_{\min}$ $=\frac{M-\lambda_{l}}{\lambda_{u}-\lambda_{l}}\lambda_{u}+\frac{\lambda_{u}-M}{\lambda_{u}-\lambda_{l}}\lambda_{l},\quad\|x^{*}_{\min}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{u}}+\frac{1}{\lambda_{l}}-\frac{M}{\lambda_{u}\lambda_{l}}.$
-This concludes the proof. $\square$
-Appendix D Other Proofs
-D.1 Proof of Proposition 3
-By Propostion 1, we have
-${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
-$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
-$\geq 1-\frac{\delta}{2}.$
-Similarly, by Definition 2, we have
-${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
-$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
-$\geq 1-\frac{\delta}{2}.$
-Combining the two inequalities above, we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta$$\square$
-D.2 Proof of Proposition 4
-Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
-$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (9)
-In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. By direct computation
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $=\left(\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle\right)+\left(\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\right)$
-$\quad+\left(\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle\right)+\left(\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right).$
-To bound the second term on the right hand side, recall by Equation (3) the POFUL action $\widetilde{x}_{t}$ satisfies
-$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
-Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
-The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (9). As a result, we have
-$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
-$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}),$
-$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}).$
-Combing all terms above, we have
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
-$\square$
-D.3 Proof of Theorem 1
-We formally prove Theorem 1 for completeness. The proof techniques are developed in previous papers [Abbasi-Yadkori et al., 2011, Agrawal and Goyal, 2013, Abeille et al., 2017].
-Throughout the proof, we condition on the event $\mathcal{A}_{T}$, which holds with probability $1-\delta$ by Proposition 3. Applying Proposition 4, we obtain
-$\mathcal{R}(T)$ $\leq\sum_{t=1}^{T}\left(\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right)\mathbb{I}\{\mathcal{A}_{t}\}$
-$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
-Recall $\|x^{\star}_{t}\|_{V^{-1}_{t}}=\alpha_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}$ and $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$, we have
-$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
-Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
-$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
-This concludes the proof.$\square$
-D.4 Proof of Proposition 5
-This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3.
-Without loss of generality, we let $M>\lambda_{2}$.
-Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$$\square$
-D.5 Proof of Theorem 3
-Define $x^{\dagger}(\theta)\coloneqq\theta/\|\theta\|$.
-The conclusion of Theorem 2 holds for $x^{\dagger}(\theta)$, i.e.
-for all $t\in[1,T]$ and all $s\leq t$, we have $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s}$ and $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}$.
-For any $\theta$, the feasible optimal action $x^{\star}_{s}(\theta)$ is the point in $\mathcal{X}_{s}$ that has the smallest distance from $x^{\dagger}(\theta)$.
-Then we have $\|x^{\dagger}(\theta)-x^{\star}_{s}(\theta)\|\leq\rho_{s}$ by the definition of the Euclidean covering radius.
-Therefore, let $\theta_{0}$ denote the $\theta$ that attains $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}$, it holds that
-$\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}=\|x^{\star}_{s}(\theta_{0})\|_{V_{s}^{-1}}$ $\geq\|x^{\dagger}(\theta_{0})\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
-$\geq\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
-$\geq\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}).$
-Similarly, we have $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$.
-As a result, it holds that $\alpha_{s}\leq(\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))/(\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))$$\square$
-Appendix E TS-MR and Greedy-MR Algorithms
+The pseudocode is presented in Algorithm 2.
 Algorithm 2 TS-MR (Greedy-MR)
@@
   end for
+By construction, the course-corrected algorithms have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$.
+Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
+Corollary 1
+TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
+Remark 4
+In practice, the choice of the threshold $\mu$ depends on the problem settings.
+In a high-risk setting where LinTS and Greedy fail with high probability, one can set a small $\mu$ so that TS-MR and Greedy-MR select more OFUL actions to guarantee the necessary amount of exploration.
+In a low-risk setting where original LinTS and Greedy algorithms work well, one can set a large $\mu$ and hence TS-MR and Greedy-MR select TS and greedy actions respectively to avoid unnecessary exploration and save the computational cost.
+7 Simulations
+In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
+7.1 Synthetic datasets
+(a) example 1
+(b) example 2
+(c) example 3
+Figure 3: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
+Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
+We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm.
+At each time $t$, we generate 100 i.i.d.
+random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$.
+This is a basic example of standard stochastic linear bandit problems without any extra structure.
+We set the threshold $\mu=8$ for TS-MR and Greedy-MR.
+We average simulation results over 100 independent runs.
+The results are summarized in Figure 2(a).
+TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
+Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
+We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$.
+At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0.
+Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks.
+This problem is equivalent to a 10-armed contextual bandit.
+We ser $\mu=12$ for TS-MR and Greedy-MR.
+Again, we average simulation results over 100 independent runs.
+The results are summarized in Figure 2(b).
+In this setting, Greedy performs suboptimally due to a lack of exploration for some arms.
+Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects large regret proxy $\widehat{\mu}_{t}$.
+Example 3. Prior mean mismatch (Hamidi and Bayati, 2020b).
+This is an example in which LinTS is shown to incur linear Bayesian regret.
+We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$.
+It is shown in Hamidi and Bayati (2020b) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$.
+Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action.
+We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit.
+We ser $\mu=12$ for TS-MR and Greedy-MR.
+We carried out 100 independent runs.
+The results are summarized in Figure 2(c).
+We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
+7.2 Real-world datasets
+We explore the performance of standard POFUL algorithms and the proposed TS-MR and Greedy-MR algorithms on real-world datasets.
+We consider three datasets for classification tasks on the OpenML platform, Cardiotocography, JapaneseVowels and Segment.
+They focus on healthcare, pattern recognition, and computer vision problems respectively.
+Setup.
+We follow the standard approach in the literature that converts classification tasks to contextual bandit problems, and then embeds it into linear bandit problems in the same way as Example 2 in Section 6.
+Specifically, we regard each class as an action so that at each time, the decision-maker assigns a class to the feature observed and receives a binary reward, namely 1 for assigning the correct class and 0 otherwise, plus a Gaussian noise.
+We plot the cumulative regret (averaged over 100 duplicates) for all algorithms.
+Figure 4 shows that for all real-world datasets: OFUL and TS-Freq perform poorly due to their conservative exploration; LinTS and Greedy are achieving empirical success even though they don’t have theoretical guarantees; TS-MR and Greedy-MR retain the desirable empirical performance of LinTS and Greedy, while enjoying the minimax optimal frequentist regret bound.
+(a) Cardiotocography dataset.
+(b) JapaneseVowels dataset.
+(c) Segment dataset.
+Figure 4: Cumulative regret of all algorithms on real-world datasets.
+8 Conclusion
+In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
+Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm.
+Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest.
+As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail.
+We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
+{APPENDICES}
+9 Proof of Lemmas for Theorem 2
+9.1 Proof of Lemma 1
+Proof.
+Let $\theta=tx$ where $x$ is any unit vector in $\mathbb{R}^{d}$ and $t\in\mathbb{R}^{+}$ is a scalar.
+Consider the equation that characterizes the intersection $\{tx:t\in\mathbb{R}^{+}\}\cap\mathcal{E}_{t}$, namely $(tx-\widehat{\theta}_{t})^{\top}V_{t}(tx-\widehat{\theta}_{t})\leq\beta_{t}^{2}$.
+Equivalently, we have $t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}\leq 0$.
+This quadratic inequality of $t$ has at least one solution if the discriminant is non-negative, i.e.
+$4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$.
+Then by direct computation,
+$\|x\|_{V_{t}}\leq\frac{\sqrt{(x^{\top}V_{t}\widehat{\theta}_{t})^{2}}}{\phi_{t}}\leq\frac{\sqrt{x^{\top}x}\sqrt{(\widehat{\theta}_{t})^{\top}V_{t}^{\top}V_{t}\widehat{\theta}_{t}}}{\phi_{t}}=\frac{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}}{\phi_{t}}.$
+Note that $\theta\in\mathcal{E}_{t}$ if and only if $\theta$ is on a ray starting from the origin and intersects $\mathcal{E}_{t}$ at one or more points.
+Namely, $\theta=tx$ for some $x$ that satisfies $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, we conclude the proof.
+9.2 Proof of Lemma 2
+Proof.
+Note that for any $\theta\in\mathcal{H}_{t}\subset\mathcal{E}_{t}$, it holds that $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$.
+Also, by the construction of $\mathcal{H}_{t}$, we have $\|\theta\|_{V_{t}}=\phi_{t}$.
+Then by direct computation, we have
+$\|x^{\star}(\theta)\|_{V_{t}}=\frac{\|\theta\|_{V_{t}}}{\|\theta\|}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
+To prove the same result for any $\theta\in\mathcal{E}_{t}$, we only need to show there exists $\theta^{\prime}\in\mathcal{H}_{t}$ such that $x^{\star}(\theta)=x^{\star}(\theta^{\prime})$.
+To see this, let $x=\theta/\|\theta\|$ and consider the intersection $\{tx:t\in\mathbb{R}^{+}\}\cap\mathcal{E}_{t}$, which is non-empty by our choice of $\theta$.
+Similar to the proof of Lemma 1, the discriminant is non-negative, i.e.
+$4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$.
+Now consider the intersection $\{tx:t\in\mathbb{R}\}\cap\partial\mathcal{E}_{t}$, where we let $\partial\mathcal{E}_{t}\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}=\beta_{t}(\delta)\right\}$ be the border of the ellisoid.
+The intersection points are characterized by the solution to
+$t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=0.$ (9)
+If $4(x^{\top}V_{t}\widehat{\theta})^{2}=4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$ and there is only one intersection point, namely $\theta$ itself, we have
+$0=t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=t^{2}\|x\|_{V_{t}}^{2}-2t\|x\|_{V_{t}}\phi_{t}+\phi_{t}^{2}=(\|tx\|_{V_{t}}-\phi_{t})^{2}.$
+Therefore, we have $\|\theta\|_{V_{t}}=\|tx\|_{V_{t}}=\phi_{t}$, i.e. $\theta\in\mathcal{H}_{t}$.
+If $4(x^{\top}V_{t}\widehat{\theta})^{2}>4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, it follows that $x^{\top}V_{t}\widehat{\theta}>\|x\|_{V_{t}}\phi_{t}$.
+This inference is valid given that $x^{\top}V_{t}\widehat{\theta}>0$, which can be verified using Equation (9) and noting that $t>0$.
+Consider the solutions to (9)
+$t_{1}=\frac{x^{\top}V_{t}\widehat{\theta}_{t}-\sqrt{(x^{\top}V_{t}\widehat{\theta})^{2}-\|x\|_{V_{t}}^{2}\phi_{t}^{2}}}{\|x\|_{V_{t}}^{2}},\quad t_{2}=\frac{x^{\top}V_{t}\widehat{\theta}_{t}+\sqrt{(x^{\top}V_{t}\widehat{\theta})^{2}-\|x\|_{V_{t}}^{2}\phi_{t}^{2}}}{\|x\|_{V_{t}}^{2}}.$
+We only need to show $\|t_{1}x\|_{V_{t}}<\phi_{t}<\|t_{2}x\|_{V_{t}}$, then by the continuity of $\|\cdot\|_{V_{t}}$ and the convexity of $\mathcal{E}_{t}$, there exists $t^{\prime}\in(t_{1},t_{2})$ such that $\|t^{\prime}x\|_{V_{t}}=\phi_{t}$.
+Then $\theta^{\prime}\coloneqq t^{\prime}x\in\mathcal{H}_{t}$ is the desired point.
+By direct computation,
+$\|t_{1}x\|_{V_{t}}=\frac{x^{\top}V_{t}\widehat{\theta}_{t}-\sqrt{(x^{\top}V_{t}\widehat{\theta})^{2}-\|x\|_{V_{t}}^{2}\phi_{t}^{2}}}{\|x\|_{V_{t}}}.$
+We only need to prove
+$x^{\top}V_{t}\widehat{\theta}_{t}-\|x\|_{V_{t}}\phi_{t}\leq\sqrt{(x^{\top}V_{t}\widehat{\theta})^{2}-\|x\|_{V_{t}}^{2}\phi_{t}^{2}}$
+Note that
+$(x^{\top}V_{t}\widehat{\theta}_{t}-\|x\|_{V_{t}}\phi_{t})^{2}$ $=(x^{\top}V_{t}\widehat{\theta})^{2}-2x^{\top}V_{t}\widehat{\theta}\phi_{t}\|x\|_{V_{t}}+\|x\|_{V_{t}}^{2}\phi_{t}^{2}$
+$\leq(x^{\top}V_{t}\widehat{\theta})^{2}-2\|x\|_{V_{t}}^{2}\phi_{t}^{2}+\|x\|_{V_{t}}^{2}\phi_{t}^{2}$
+$=(x^{\top}V_{t}\widehat{\theta})^{2}-\|x\|_{V_{t}}^{2}\phi_{t}^{2}$
+where we have used the fact that $x^{\top}V_{t}\widehat{\theta}>\|x\|_{V_{t}}\phi_{t}$ in the ineuqlity.
+Taking square root for both sides yields the desired result, and hence $\|t_{1}x\|_{V_{t}}<\phi_{t}$.
+Similarly, one can show $\phi_{t}<\|t_{2}x\|_{V_{t}}$.
+This concludes the proof.
+9.3 Proof of Lemma 3
+Proof.
+Let $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{V}}\}$ be the set of distinct eigenvalues of $V$.
+By the spectral theorem, $V$ can be decomposed as $V=U\Lambda U^{\top}$, where the columns of $U$ consist of orthonormal eigenvectors of $V$, denoted by $\{u_{11},\dots,u_{1n_{1}},u_{21},\dots,u_{2n_{2}},\dots,u_{N_{V}1},\dots,u_{N_{V}n_{N_{V}}}\}$, where $n_{1},\dots,n_{N_{V}}$ are the algebraic multiplicity of the eigenvalues respectively.
+Since $V$ is a symmetric matrix, the eigenvectors for a basis of $\mathbb{R}^{d}$ and we have $\sum_{i=1}^{N_{V}}n_{i}=d$.
+We can write $x$ as a linear combination $x=\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}u_{ij}$, where $\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}^{2}=1$.
+Define $a_{i}\coloneqq\sum_{j=1}^{n_{i}}w_{ij}^{2}$, by direct computation, we have $\sum_{i=1}^{N_{V}}a_{i}=1$ and $\|x\|_{V}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}a_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}^{-1}a_{i}$
+Now we study the range of $\|x\|_{V^{-1}}^{2}$ when $\|x\|_{V}^{2}$ is bounded as $m\leq\|x\|_{V}^{2}\leq M$. First, let’s focus on maximizing $\|x\|_{V^{-1}}^{2}$, it suffices to solve the LP problem
+maximize $\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}^{-1}\quad\text{s.t.}\quad\forall i,a_{i}\geq 0,\ \sum_{i=1}^{N_{V}}a_{i}=1,\ \sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\ \sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M.$
+The Lagrangian is given by
+$L=\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}^{-1}+\mu(1-\sum_{i=1}^{N_{V}}a_{i})+\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)+\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})+\sum_{i=1}^{N_{V}}\kappa_{i}a_{i}.$
+The KKT conditions are given by
+$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{N_{V}}a_{i}=1,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$ (10)
+To satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices.
+Hence, $a_{i}$ can only be non-zero for at most two distinct eigenvalues, denoted by $\lambda_{i}$ and $\lambda_{j}$, where $i<j$ and $\lambda_{i}>\lambda_{j}$.
+Namely, the solution to (10) lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{i}$ and $\lambda_{j}$.
+Let $y=\|x\|_{V}^{2}$, we have $a_{i}\lambda_{i}+a_{j}\lambda_{j}=y$ and $a_{i}\lambda_{i}^{-1}+a_{j}\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
+Note that $\sum_{i=1}^{N_{V}}a_{i}=a_{i}+a_{j}=1$, by direct computation, the closed form of $\|x\|_{V^{-1}}^{2}$ is given by $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$.
+Clearly, we have
+$\begin{cases}\frac{\partial f}{\partial y}&=-\frac{1}{\lambda_{i}\lambda_{j}}<0,\\ \frac{\partial f}{\partial\lambda_{i}}&=(\frac{y}{\lambda_{j}}-1)\frac{1}{\lambda_{i}^{2}}>0,\\ \frac{\partial f}{\partial\lambda_{j}}&=(\frac{y}{\lambda_{i}}-1)\frac{1}{\lambda_{j}^{2}}<0.\end{cases}$
+Then the maximum of $\|x\|_{V^{-1}}^{2}$ is obtained when $\lambda_{i}=\lambda_{1}$, $\lambda_{j}=\lambda_{N_{V}}$, $y=m$.
+Therefore, the solution to the LP problem is any unit vector $x^{\star}_{\max}$ that lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{1}$ and $\lambda_{N_{V}}$.
+Moreover, we have
+$\|x^{\star}_{\max}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{N_{V}}}-\frac{m}{\lambda_{1}\lambda_{N_{V}}}.$
+Similarly, by investigating the KKT conditions for the LP problem that minimize $\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}^{-1}$, the minimum of $\|x\|_{V^{-1}}^{2}$ is obtained when $\lambda_{i}=\lambda_{k}$, $\lambda_{j}=\lambda_{k+1}$, $y=M$, where $k$ is such that $\lambda_{k}=\max_{i\in[N_{v}]}\{\lambda_{i}\geq M\}$, and hence $\lambda_{k+1}=\min_{i\in[N_{v}]}\{\lambda_{i}<M\}$.
+The solution vector is any unit vector $x^{\star}_{\min}$ that lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{k}$ and $\lambda_{k+1}$, and we have
+$\|x^{\star}_{\min}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}.$
+This concludes the proof.
+10 Other Proofs
+10.1 Proof of Proposition 3
+Proof.
+By Propostion 1, we have
+${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
+$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
+$\geq 1-\frac{\delta}{2}.$
+Similarly, by Definition 2, we have
+${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
+$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
+$\geq 1-\frac{\delta}{2}.$
+Combining the two inequalities above, we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta$. ∎
+10.2 Proof of Proposition 4
+Proof.
+Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
+$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (11)
+In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. We have
+$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $=\left(\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle\right)+\left(\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\right)$
+$\quad+\left(\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle\right)+\left(\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right).$
+To bound the second term on the right hand side, recall by Equation (3) the POFUL action $\widetilde{x}_{t}$ satisfies
+$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
+Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
+The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (11). As a result, we have
+$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
+$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t},$
+$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
+Combing all terms above, we have
+$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
+10.3 Proof of Theorem 1
+Proof.
+We formally prove Theorem 1 for completeness. The proof techniques are developed in previous papers (Abbasi-Yadkori et al., 2011; Agrawal and Goyal, 2013; Abeille et al., 2017).
+Throughout the proof, we condition on the event $\mathcal{A}_{T}$, which holds with probability $1-\delta$ by Proposition 3. Applying Proposition 4, we obtain
+$\mathcal{R}(T)$ $\leq\sum_{t=1}^{T}\left(\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right)\mathbb{I}\{\mathcal{A}_{t}\}$
+$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
+Recall $\|x^{\star}_{t}\|_{V^{-1}_{t}}=\alpha_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}$ and $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$, we have
+$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
+Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
+$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
+This concludes the proof. ∎
+10.4 Proof of Proposition 5
+Proof.
+This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3.
+Without loss of generality, we let $M>\lambda_{2}$.
+Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$.
+∎
+11 Empirical Validation of Case 2 from Section 5.2
+To empirically confirm that Case 2, as discussed in Section 5.2, is not merely theoretical, we examine the condition outlined in Proposition 5.
+Our focus is on the ratio $\zeta_{t}\coloneqq\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}/\|\widehat{\theta}_{t}\|^{2}}{\lambda_{1}(V_{t})}$.
+Notably, $\|\widehat{\theta}_{t}\|_{V_{t}}^{2}$ tends to approximate $\lambda_{1}\|\widehat{\theta}_{t}\|^{2}$ when $\widehat{\theta}_{t}$ is proximate to the top eigenspace of $V_{t}$.
+Consequently, this ratio serves as a proxy of the alignment of $\widehat{\theta}_{t}$ with the top eigenspace of $V_{t}$.
+Specifically, a $\zeta_{t}$ value of 1 signifies that $\widehat{\theta}_{t}$ is within the top eigenspace.
+We re-evaluate Example 1 from Section 7, which epitomizes the general linear bandit problem in its standard form. The empirical sequence ${\zeta_{t}}$ is depicted in Figure 5.
+Figure 5: Evolution of the alignment proxy $\zeta_{t}$ in Example 1.
+Figure 5 reveals that for every bandit algorithm, the value of $\zeta_{t}$ is notably high (exceeding 0.9) at the early stages of implementation and gradually converges towards 1.
+This observation validates the behavior outlined in Case 2.
+Furthermore, the consistency of this phenomenon across all examined bandit algorithms suggests that such a tendency is likely a characteristic of the online nature of these algorithms.
+It is important to note that the initial sharp decline observed is attributed to the behavior of the regularized least squares estimator in over-parameterized scenarios, when $t$ is less than the dimension $d$.

### v5

@@
-
+Geometry-Aware Bandit Algorithms
 […]
 […]
+Geometry-Aware Bandit Algorithms
+[…]
+Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits
+[…]
+[…]
+[…]
+This paper is motivated by recent research in the $d$-dimensional stochastic linear bandit literature, which has revealed an unsettling discrepancy: algorithms like Thompson sampling and Greedy demonstrate promising empirical performance, yet this contrasts with their pessimistic theoretical regret bounds.
+The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances.
+To address this, we propose a new data-driven technique that tracks the geometric properties of the uncertainty ellipsoid around the main problem parameter.
+This methodology enables us to formulate a data-driven frequentist regret bound, which incorporates the geometric information, for a broad class of base algorithms, including Greedy, OFUL, and Thompson sampling.
+This result allows us to identify and “course-correct” problem instances in which the base algorithms perform poorly.
+The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$ for a $T$-period decision-making scenario, effectively maintaining the desirable attributes of the base algorithms, including their empirical efficacy.
+We present simulation results to validate our findings using synthetic and real data.
+[…]
+Linear bandit, Thompson sampling, Greedy, Data-driven exploration
 1 Introduction
-Multi-armed bandits (MABs) have garnered significant attention as they provide well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems.
-In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards.
-This paper focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases.
-In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with mean equal to the inner product of $\theta^{\star}$ and the chosen action.
+Multi-armed bandits (MABs) provide a framework for studying the exploration-exploitation trade-off in sequential decision-making, where a decision-maker selects actions and observes uncertain rewards.
+This extends to contextual bandits with features or covariates, as shown in numerous applications (Langford and Zhang, 2008; Li et al., 2010; Tewari and Murphy, 2017; Zhou et al., 2020; Villar et al., 2015; Bastani and Bayati, 2020; Cohen et al., 2020).
+This paper focuses on a well-studied class of models that captures both MABs and contextual bandits as special cases while being amenable to theoretical analysis: the stochastic linear bandit (LB) problem.
+In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with a mean equal to the inner product of $\theta^{\star}$ and the chosen action.
 The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time.
@@
 In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards.
-Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010).
+Various algorithms based on the optimism principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010).
 Another popular strategy is Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions.
 Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
-In the linear bandit setting, two types of regret are considered.
-The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$.
-Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al.
-(2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to polylogarithmic factors.
-On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of the noise and the algorithm.
-In this frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al.
-(2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound.
-However, for a frequentist variant of LinTS referred to as TS-Freq, which inflates the posterior distribution, Agrawal and Goyal (2013) and Abeille et al.
-(2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret.
-This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$.
-Recent work by Hamidi and Bayati (2020b) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved.
-For the Greedy algorithm in linear bandit problems, no general theoretical guarantee exists.
-Moreover, it is known that both LinTS and Greedy might perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
-Despite the theoretical gaps in performance, LinTS has shown strong empirical performance (Russo et al., 2018), indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently.
-Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021).
-Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al.
-(2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency.
-This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify, in a data-driven fashion, problematic instances where LinTS and Greedy could potentially fail and apply a “course-correction” to ensure competitive frequentist regret bounds?
-And can this be achieved without compromising their impressive empirical performance and computational efficiency?
+In the linear bandit setting, two regret types are considered.
+Bayesian regret treats parameter $\theta^{\star}$ as a random variable with a prior distribution, averaging regret over noise, algorithm randomness, and parameter randomness, measuring expected performance across parameter realizations.
+Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the Thompson Sampling (TS) heuristic, referred to as LinTS, matching the minimax optimal bound by Dani et al.
+(2008).
+Here, $\widetilde{\mathcal{O}}$ denotes asymptotic order up to polylogarithmic factors.
+Frequentist regret assumes fixed unknown parameter $\theta^{\star}$, averaging only over noise and algorithm randomness.
+The OFUL algorithm (Abbasi-Yadkori et al., 2011) achieves an optimal $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound.
+However, TS-Freq, a frequentist LinTS variant with inflated posterior variance, only achieves $\widetilde{\mathcal{O}}(d\sqrt{dT})$ (Agrawal and Goyal, 2013; Abeille et al., 2017), suboptimal by factor $\sqrt{d}$.
+Hamidi and Bayati (2020b) confirms this inflation is necessary and LinTS’s frequentist regret cannot be improved.
+The Greedy algorithm lacks theoretical guarantees for linear bandit problems (Lattimore and Szepesvari, 2017), suggesting both LinTS and Greedy may perform suboptimally.
+Despite the theoretical gaps, LinTS demonstrates strong empirical performance (Russo et al., 2018), suggesting posterior distribution inflation may be unnecessary in most scenarios.
+Similarly, the Greedy algorithm performs well in typical cases (Bietti et al., 2021).
+While optimism-based algorithms are computationally expensive (generally NP-hard (Dani et al., 2008; Russo and Van Roy, 2014; Agrawal, 2019)), LinTS and Greedy maintain computational efficiency.
+This disparity between theoretical, computational, and empirical performance prompts two questions: Can we identify problematic instances for LinTS and Greedy in a data-driven way and apply ”course-correction” to ensure competitive frequentist regret bounds?
+Can this be achieved while preserving their empirical performance and computational efficiency?
 In this paper, we provide positive answers to both questions.
 Specifically, we make the following contributions.
+1.
 We develop a real-time geometric analysis technique for the $d$-dimensional confidence ellipsoid surrounding $\theta^{\star}$.
@@
 Consequently, this facilitates a more precise “course-correction”.
-  • We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, instance-dependent frequentist regret bound for them.
+2.
+We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, data-driven frequentist regret bound for them.
 This bound is efficiently computable using data observed from previous decision epochs.
+3.
 We introduce course-corrected variants of LinTS and Greedy that achieve minimax optimal frequentist regret.
 These adaptations maintain most of the desirable characteristics of the original algorithms.
-The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid.
-This is in contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their focus on summary statistics.
 1.1 Other Related Literature
-We note that Valko et al.
-(2014); Kocák et al.
-(2014, 2020); Kocák and Garivier (2020) consider the spectral of graphs and study Spectral Bandits.
-In their work, the rewards of arms are represented as smooth functions on a graph, allowing the exploitation of a low-rank structure for improved algorithmic performance and a regret guarantee that is independent of the original number of actions.
-In contrast, our approach leverages the spectral properties of the sample covariance of actions, which is distinct from graph spectral analysis.
-Additionally, our research addresses the broader context of stochastic linear bandit problems without presupposing any low-rank structure.
-The concepts in this paper bear similarities to existing research that investigates various exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018), as well as recent studies by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a) that focus on using observed data for reducing exploration.
-Nonetheless, our methodology and the specific use of data markedly differ.
-For example, Bastani et al.
-(2021) emphasizes the minimum eigenvalue of the sample covariance matrix, essentially a single-parameter summary of the observed data.
-On the other hand, Hamidi and Bayati (2020a) uses information from the one-dimensional confidence intervals surrounding the reward values.
-A more closely related paper, Hamidi and Bayati (2020b), uses spectral information to enhance the performance of TS in LB problems.
-They look at a single summary statistics of the confidence set that they call thinness coefficient, to decide if the posterior of TS should be inflated.
-In contrast, with these studies, our approach capitalizes on the full geometric details of the $d$-dimensional confidence ellipsoid, utilizing richer geometric information.
-1.2 Organization
-The structure of this paper is as follows: Section 2 introduces the notations and problem formulation.
-In Section 3-4, we present the POFUL family of algorithms and analyze their frequentist regret.
-Sections 5-6 detail our data-driven approach and meta-algorithm for course-correction.
-Section 7 discusses simulations using synthetic and real-world datasets.
-The paper concludes in Section 8 with a summary of findings.
-All technical proofs, further discussions, and supplementary materials are provided in the appendices.
-endcomment
+Our work is closely related to three main research streams: methodological foundations of linear bandits, bandit algorithms utilizing spectral properties, and data-driven exploration techniques.
+While these works share some similarities with our approach, we highlight the key differences and the unique aspects of our methodology.
+From a methodological perspective, our regret analysis builds upon the foundations laid by Abbasi-Yadkori et al.
+(2011), Agrawal and Goyal (2013), and Abeille et al.
+(2017).
+However, a key distinguishing factor is that our approach does not rely on optimistic samples, which is a departure from previous methods.
+This means that the algorithms we study do not always choose actions that are expected to perform better than the true optimal action.
+By allowing non-optimistic samples, we avoid the need to inflate the posterior distribution, a requirement in the works of Agrawal and Goyal (2013) and Abeille et al.
+(2017).
+Our use of spectral information in bandit algorithms bears some resemblance to the study of Spectral Bandits (Valko et al., 2014; Kocák et al., 2014, 2020; Kocák and Garivier, 2020).
+These works represent arm rewards as smooth functions on a graph, leveraging low-rank structures to improve algorithmic performance and obtain regret guarantees independent of the number of actions.
+In contrast, our approach exploits the spectral properties of the action covariance matrix, which is distinct from graph spectral analysis.
+Moreover, our research tackles the broader context of stochastic linear bandits without assuming any low-rank structure.
+Our work also shares conceptual similarities with research on exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018) and data-driven exploration reduction (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a).
+However, our methodology and data utilization differ significantly.
+For instance, Bastani et al.
+(2021) focuses on the minimum eigenvalue of the covariance matrix, a single-parameter summary of the observed data, while Hamidi and Bayati (2020a) uses information from one-dimensional reward confidence intervals.
+The work of Hamidi and Bayati (2020b) is more closely related to ours, as it employs spectral information to improve the performance of Thompson Sampling in linear bandits.
+They use a single summary statistic called the thinness coefficient to decide whether to inflate the posterior.
+In contrast, our approach leverages the full geometric details of the $d$-dimensional confidence ellipsoid, harnessing richer geometric information.
 2 Setup and Preliminaries
@@
 […]
-Problem Formulation.
+Problem formulation and assumptions.
 We consider the stochastic linear bandit problem.
@@
 At each time $t\in[T]$, a policy $\pi$ selects action $x_{t}$ from a set of action $\mathcal{X}_{t}\subset\mathbb{R}^{d}$ according to the past observations and receives a reward $r_{t}=\langle{x_{t}},{\theta^{\star}}\rangle+\varepsilon_{t}$, where $\varepsilon_{t}$ is mean-zero noise with a distribution specified in Assumption 3 below.
-We measure the performance of $\pi$ with the cumulative expected regret $\mathcal{R}(T)=\sum_{t=1}^{T}\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x_{t}},{\theta^{\star}}\rangle,$ where $x^{\star}_{t}$ is the best action at time $t$, i.e., $x^{\star}_{t}=\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$.
-Let $\mathcal{F}_{t}$ be a $\sigma$-algebra generated by the history $(x_{1},r_{1},\ldots,x_{t},r_{t})$ and the prior knowledge, $\mathcal{F}_{0}$.
+We measure the performance of $\pi$ with the cumulative expected regret $\mathcal{R}(T)=\sum_{t=1}^{T}\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x_{t}},{\theta^{\star}}\rangle\,,$ where $x^{\star}_{t}$ is the best action at time $t$, i.e., $x^{\star}_{t}=\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle\,.$ Let $\mathcal{F}_{t}$ be a $\sigma$-algebra generated by the history $(x_{1},r_{1},\ldots,x_{t},r_{t})$ and the prior knowledge, $\mathcal{F}_{0}$.
 Therefore, $\{\mathcal{F}_{t}\}_{t\geq 0}$ forms a filteration such that each $\mathcal{F}_{t}$ encodes all the information up to the end of period $t$.
-2.1 Assumptions
 We make the following assumptions that are standard in the relevant literature.
@@
 The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$.
-Assumption 3 (Noise)
+Assumption 3 (Subgaussian reward noise)
 The noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ is conditionally mean-zero and $R$-subgaussian, where $R$ is known.
@@
 2.1 Regularized Least Square and Confidence Ellipsoid
-Here, we review useful frequentist tools developed in Abbasi-Yadkori et al. (2011).
-Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards.
-In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator.
-Fix a regularization parameter $\lambda$, the sample covariance matrix, and the Regularized Least Square (RLS) estimate are respectively given by
-$V_{t}=\lambda I+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
-The following proposition in Abbasi-Yadkori et al. (2011) states that the RLS estimate $\widehat{\theta}_{t}$ concentrates around $\theta^{\star}$ properly.
+In this subsection, we review the useful frequentist tools developed by Abbasi-Yadkori et al. (2011) for estimating the unknown parameter $\theta^{*}$ in linear bandit (LB) problems.
+Consider an arbitrary sequence of actions $(x_{1},\dots,x_{t})$ and their corresponding rewards $(r_{1},\dots,r_{t})$.
+In LB problems, the parameter $\theta^{*}$ is typically estimated using the regularized least squares (RLS) estimator.
+Let $\lambda_{\text{reg}}$ be a fixed regularization parameter.
+The sample covariance matrix $V_{t}$ and the RLS estimate $\widehat{\theta}_{t}$ are defined as follows:
+$V_{t}=\lambda_{\text{reg}}I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
+The following proposition from Abbasi-Yadkori et al. (2011) establishes that the RLS estimate $\widehat{\theta}_{t}$ concentrates around the true parameter $\theta^{*}$ with high probability.
 Proposition 1 (Theorem 2 in Abbasi-Yadkori et al. (2011))
-Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that $\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)$ and $\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$ for all $x\in\mathbb{R}^{d}$, where
-$\beta_{t}^{RLS}(\delta)=R\sqrt{2\log\frac{(\lambda+t)^{d/2}\lambda^{-d/2}}{\delta}}+\sqrt{\lambda}S.$ (2)
-Hereafter, we might omit the dependence of $\beta_{t}$ on $\delta$ if there is no ambiguity. Propostion 1 enables us to construct the following sequence of confidence ellipsoids.
+Let $\delta\in(0,1)$ be a fixed confidence level. Then, with probability at least $1-\delta$, it holds for all $x\in\mathbb{R}^{d}$ that
+$\|\widehat{\theta}_{t}-\theta^{\star}\|_{V_{t}}\leq\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS},\quad|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle|\leq\|x\|_{V_{t}^{-1}}\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}$
+where the confidence bound $\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}$ is defined as
+$\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}=R\sqrt{2\log(\lambda_{\text{reg}}+t)^{d/2}\lambda_{\text{reg}}^{-d/2}{\delta}^{-1}}+\sqrt{\lambda_{\text{reg}}}S.$ (2)
+Proposition 1 enables us to construct the following sequence of confidence ellipsoids.
 Definition 1
-Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as $\mathcal{E}_{t}^{RLS}(\delta)=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)\}$.
-The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret.
-Essentially, this proposition suggests that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
+Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as
+$\mathcal{E}_{t,\delta,\lambda_{\text{reg}}}^{RLS}=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}\}\,.$
+The next proposition, known as the elliptical potential lemma, plays a central role in bounding the regret.
+This proposition provides the key element in the work of Abbasi-Yadkori et al.
+(2011), showing that the cumulative prediction error incurred by the action sequence used to estimate $\theta^{*}$ is small.
 Proposition 2 (Lemma 11 in Abbasi-Yadkori et al. (2011))
-If $\lambda>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda I)}\leq 2d\log\left(1+\frac{t}{\lambda}\right).$
+If $\lambda_{\text{reg}}>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda_{\text{reg}}I)}\leq 2d\log(1+\frac{t}{\lambda_{\text{reg}}})\,.$
 3 POFUL Algorithms
-In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). It allows us to analyze the frequentist regret of common algorithms in a unified way.
+In this section, we introduce POFUL (Pivot OFUL), a generalized framework of OFUL. This framework enables a unified analysis of frequentist regret for common algorithms.
 At a high level, POFUL is designed to encompass the exploration mechanism of OFUL and LinTS.
@@
 The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability.
-This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
+This is formalized in the next definition.
 Definition 2
 Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$.
-Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$,
-$\mathbb{P}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]\geq 1-\delta^{\prime},$
-where we define the pivot ellipsoid as $\mathcal{E}_{t}^{PVT}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t}(\delta)\right\}$.
+Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$, $\mathbb{P}[\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}]\geq 1-\delta^{\prime}$, where we define the “pivot ellipsoid” as $\mathcal{E}_{t,\delta,\lambda_{\text{reg}}}^{PVT}\coloneqq\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t,\delta,\lambda_{\text{reg}}}\}$.
 At each time $t$, POFUL chooses the action that maximizes the optimistic reward
-$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$ (3)
-A pseudocode representation of POFUL is provided in Algorithm 1.
-Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid.
-We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$.
-By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
+$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\,,$ (3)
+as shown in a pseudocode representation in Algorithm 1 and illustrated in Figure 1(a).
+Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid.
+We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$.
 Algorithm 1 POFUL
-0:  $T$, $\delta$, $\lambda$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
-  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
+0:  $T$, $\delta$, $\lambda_{\text{reg}}$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
+  Initialize $V_{0}\leftarrow\lambda_{\text{reg}}I_{d}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
   for t = 0, 1, …, T  do
      Sample a feasible pivot $\widetilde{\theta}_{t}$ with respect to $\iota_{t}$ according to Definition 2
-     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$
+     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$
      Observe reward $r_{t}$
@@
   end for
-OFUL
-For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$ Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
-LinTS and TS-Freq
+[…]
+[…]
+Figure 1: (a) POFUL algorithms illustration for general $\iota_{t}$ and $\tau_{t}$. (b) Special cases: OFUL ($\iota_{t}=0$, $\tau_{t}=1$), TS ($\tau_{t}=0$), and Greedy ($\iota_{t}=\tau_{t}=0$).
+We demonstrate that POFUL encompasses OFUL, LinTS, TS-Freq, and Greedy as special cases, as illustrated in Figure 1(b).
+Example 1 (OFUL)
+For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$.
+Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
+Before describing how TS can be derived as an instance of POFUL, we introduce a definition.
+Definition 3
+Let $\delta\in(0,1)$. We define $\mathcal{D}^{SA}(\delta)$ as a distribution satisfying $\mathbb{P}_{\eta\sim\mathcal{D}^{SA}(\delta)}\left[\|\eta\|\leq 1\right]\geq 1-\delta$.
+Example 2 (TS)
 Linear Thompson Sampling (LinTS) algorithm is a generic randomized algorithm that samples from a distribution constructed from the RLS estimate at each step.
-At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, and $\iota^{TS}_{t}$ is an inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability.
-Definition 3
-Fix any $\delta\in(0,1)$.
-$\mathcal{D}^{SA}(\delta)$ is a distribution such that $\mathbb{P}_{\theta\sim\mathcal{D}^{SA}(\delta)}\left[\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq 1\right]\geq 1-\delta.$
-LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$.
-We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm.
-On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al.
+At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-1/2}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, $\iota^{TS}_{t}$ is inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability.
+LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-1/2}\eta_{t}$.
+Setting $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm, while setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al.
 (2017), namely TS-Freq.
 This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
-Greedy
-Greedy is a special case of POFUL where $\iota_{t}=\tau_{t}=0$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$ for all $t\in[T]$.
+Example 3 (Greedy)
+Greedy is a special case of POFUL with $\iota_{t}=\tau_{t}=0$, $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, $\forall t$.
 4 Frequentist Regret Analysis of POFUL
-In this section, we present the frequentist regret analysis of POFUL algorithms. We first introduce useful concentration events that hold with high-probability.
+In this section, we present the frequentist regret analysis of POFUL algorithms. Proofs appear in Appendix 11. We first introduce high-probability concentration events.
 Definition 4
-Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. We define $\beta^{PVT}_{t}(\delta^{\prime})\coloneqq\iota_{t}(\delta)\beta^{RLS}_{t}(\delta^{\prime})$ and
-$\ \widehat{\mathcal{A}}_{t}$$\coloneqq\left\{\forall s\leq t:\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\left\{\forall s\leq t:\left\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\right\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})\right\}.$
-We also define $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$
+Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$.
+We define $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, $\widehat{\mathcal{A}}_{t}\coloneqq\{\forall s\leq t:\|\widehat{\theta}_{t}-\theta^{\star}\|_{V_{t}}\leq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\{\forall s\leq t:\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\}$, and $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$.
 Proposition 3
 Under Assumptions 1, 2 and 3 , we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta.$
-In the following, we decompose the regret of POFUL into three terms, $\mathcal{R}^{PE}(T)$, $\mathcal{R}^{PVT}(T)$ and $\mathcal{R}^{RLS}(T)$, and sketch how they are bounded separately.
-Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the benchmark action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
-$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
-$\eqqcolon\mathcal{R}^{PE}(T)+\mathcal{R}^{PVT}(T)+\mathcal{R}^{RLS}(T).$
-$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator.
-Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate.
-By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$.
-Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$.
-$\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$.
-Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$.
-As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$.
-Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound.
-For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$.
-However, later we see this could be problematic for bounding $\mathcal{R}^{PE}(T)$.
-$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term.
-In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$.
-In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$.
-For OFUL, this term is bounded properly with high probability by the construction of OFUL actions.
-For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al.
-(2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$.
-The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$.
-In Abeille et al.
-(2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e.
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ with a constant probability.
-For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures.
-The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy that cannot be analyzed by the optimism-based methods.
-In fact, non-optimistic samples might be able to control the pessimism term.
-Our method does not force choosing optimistic samples and hence there is no need to inflate the posterior.
-As a result, it does not introduce a $\sqrt{d}$-gap in the regret.
-Instead, we introduce a measurement of the “quality” of POFUL actions with respect to the regret, and develop a computable upper bound for this quality using geometric approaches.
-The main merits are an instance-dependent regret bound for POFUL that matches the minimax optimal regret in some settings, and variants of LinTS and Greedy with provable frequentist regret bound that is minimax optimal up to only a constant factor.
-4.1 An Instance-Dependent Regret Bound for POFUL
-In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$.
-The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix 10.2.
+4.1 An Data-Driven Regret Bound for POFUL
+In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL.
 Proposition 4
-Suppose ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
-On the right-hand side, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret.
-To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$-norm.
+Suppose ${\theta}^{\star}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$, it holds that
+$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$ (4)
+Note that this upper bound is different from what’s used in the optimism-based methods (Abbasi-Yadkori et al., 2011; Agrawal and Goyal, 2013; Abeille et al., 2017), we reproduce their upper bound and discuss the relationship of our method and theirs in Appendix 12.
+On the right-hand side of (4), since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}$ and get an upperbound of the regret.
+To address this, the key point to connect $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ with $V_{t}^{-1}$-norm.
 This motivates the following definition.
@@
 We also define the (instantaneous) regret proxy at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
-Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$.
+Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$.
 Hence, $\alpha_{t}$ serves as the ratio of uncertainty degrees of the reward obtained by the optimal action $x^{\star}_{t}$ and the chosen action $\widetilde{x}_{t}$.
 The intuition behind the definition for $\mu_{t}$ is constructing a regret upper bound similar to that of OFUL.
-Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$.
+Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$.
 In this sense, $\mu_{t}$ is a proxy of the instantaneous regret incurred by POFUL at time $t$.
-Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}\equiv 1$ for OFUL (recall $\iota_{t}\equiv 0$ and $\tau_{t}\equiv 1$ for OFUL).
-The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix 10.3.
-It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
+Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}=1$ for all $t\in[T]$ for OFUL (recall that in OFUL, $\iota_{t}=0$ and $\tau_{t}=1$ for all $t\in[T]$).
+The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$. It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
 Theorem 1 (Oracle frequentist regret bound for POFUL)
-Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a frequentist regret of
-$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda}\right)}\beta_{T}^{RLS}(\delta^{\prime}).$ (4)
+Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a regret of
+$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda_{\text{reg}}}\right)}\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\,.$ (5)
 Remark 1
@@
 In general, they cannot be calculated by the decision-maker.
-Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy.
-Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL, and could be utilized later for the course correction.
-This is indeed our aim in the next section and we achieve it using geometric arguments.
+Nevertheless, note that $\iota_{t}$ and $\tau_{t}$ are chosen by the decision-maker, when we have computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$, using $\mu_{t}^{2}\leq 2\alpha_{t}^{2}(1+\iota_{t}-\tau_{t})^{2}+2(1+\iota_{t}+\tau_{t})^{2}$, we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well.
+Consequently, Theorem 1 instantly turns into a data-driven regret bound for POFUL and could be utilized later for course correction, which will be the aim of the next section.
+When we additionally know that $1+\iota_{t}-\tau_{t}$ is non-negative, we would use the equality $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ directly for the bound.
 Remark 2
-A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al.
-(2017), where they indicate their proof of regret bound for LinTS can be reproduced if LinTS samples $\widetilde{\theta}_{t}$ satisfying $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$.
-They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation.
-Our work answers this question positively in the next section by investigating the property of (the reciprocal of their) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
+In the Discussion section of Abeille et al.
+(2017), the authors introduce a concept similar to the reciprocal of our $\alpha_{t}$.
+They suggest that the necessity of proving LinTS samples are optimistic could be bypassed if for some $\alpha>0$ LinTS samples $\widetilde{\theta}_{t}$ such that $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability, where $x^{\star}(\widetilde{\theta}_{t})$ and $x^{\star}(\theta^{\star}_{t})$ represent the optimal actions corresponding to $\widetilde{\theta}_{t}$ and $\theta^{\star}_{t}$, respectively.
+They pose this as an open question regarding the possibility of relaxing the requirement of inflating the posterior.
+In the following section, we provide a positive answer to this question by studying the reciprocal of their $\alpha$ using geometric arguments.
+This investigation offers an explanation for the empirical success of LinTS without the need for posterior inflation.
 5 A Data-Driven Approach
@@
 Note that $\iota_{t}$ and $\tau_{t}$ are parameters of POFUL that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$, which turns into an upper bound $\widehat{\mu}_{t}$ for the regret proxy $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
-Typically, we can check $\widehat{\alpha}_{t}$ is no less than one because we might have $\widetilde{x}_{t}=x_{t}^{\star}$.
-As a result, it holds that $\widehat{\mu}_{t}=(1+\widehat{\alpha}_{t})(1+\iota_{t})+(1-\widehat{\alpha}_{t})\tau_{t}\geq(1+\widehat{\alpha}_{t})(1+\iota_{t})$, i.e., non-zero $\tau_{t}$ gives a larger regret proxy.
-Given this observation, we’ll focus on scenarios where $\tau_{t}\equiv 0$.
-This avoids the large regret proxies and computational inefficiencies seen in OFUL when determining the action $\widetilde{x}_{t}$.
-Such scenarios include LinTS and its variants like TS-Freq as well as Greedy - all of which are standard algorithms that still lack theoretical regret guarantee results.
-In the following, we construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for both discrete and continuous action scenarios.
-5.1 Discrete Action Space
-We start with the discrete action space scenario to illustrate our method.
-It turns out that upper bounding $\alpha_{t}$ is done by simply comparing the confidence interval lengths of all potentially optimal actions.
-To see this, recall for an arbitrary action $x\in\mathcal{X}_{t}$, by Proposition 1, with high probability its expected reward is bounded as
-$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\eqqcolon U_{t}(x).$
-The confidence interval length is proportional to $\|x\|_{V_{t}^{-1}}$.
-Therefore, to characterize the range of $\|x_{t}^{\star}\|_{V_{t}^{-1}}$, we only need to identify the set of potentially optimal actions and calculate the minimal $V_{t}^{-1}$-norm among all actions in that set.
-Note that for an action $x\in\mathcal{X}_{t}$ to remain potentially optimal, it cannot be dominated by another action.
-I.e., we require $U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}L_{t}(y)$.
-Therefore, define
-$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\},$
-the set of potentially optimal actions is given by $C_{t}(\beta_{t}^{RLS}(\delta^{\prime}))$.
-Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))$, where we let $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$.
-The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
-By comparing the range of $V_{t}^{-1}$-norm for both sets, an upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$ is given by
-$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (5)
-5.2 Continuous action space.
-When it comes to the generic settings where the number of actions is large or when the set of actions is a continuous set in $\mathbb{R}^{d}$, the upper bound above is inefficient or even infeasible to compute.
-In this section, we tackle this issue.
+We focus on scenarios where $\tau_{t}=0$ for all $t\in[T]$.
+These include LinTS and variants such as TS-Freq, as well as Greedy - standard algorithms still lacking theoretical regret guarantees.
+Below, we construct upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for the continuous-action scenario.
+Bounds for discrete-action scenarios appear in Appendix 13.
+5.1 Continuous action space
 Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids, providing upper bounds that can be computed efficiently.
-For the sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for this scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al.
+For the sake of a better illustration, we consider $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$ for this scenario, where ${\mathcal{S}}_{d-1}=\left\{x\in\mathbb{R}^{d}:\|x\|=1\right\}$ is the unit hypersphere in $\mathbb{R}^{d}$.
+This is a standard example of continuous action space, and is the same as the setting considered in Abeille et al.
 (2017).
@@
 This is because we don’t have a closed-form solution for the set of potentially optimal actions.
-In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$.
-To upper bound $\alpha_{t}$, a direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$.
+In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes the form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$.
+To upper bound $\alpha_{t}$, we consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and $\mathcal{E}^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}$.
 Specifically, we have
-$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (6)
-As is illustrated in Figure 1.
-The set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
+$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (6)
+As is illustrated in Figure 2, the set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
 It’s hard to get a closed-form expression for $\mathcal{C}_{t}$, so we cannot directly calculate the range of $V_{t}^{-1}$-norm of actions in $\mathcal{C}_{t}$.
-Nevertheless, we can approximate the range by investigating the geometric properties of the ellipsoids.
-The intuition here is that, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$.
-All actions within $\mathcal{C}_{t}$ point to similar directions and thus have similar $V_{t}$-norm.
-(Note that for a unit vector $x$, we have that $\|x\|_{V_{t}}$ can be written as a weighted summation of the eigenvalues of $V_{t}$, where the weights are determined by the direction of $x$.) Hence we can bound the range of $V_{t}$-norm by investigating the concentration of $\mathcal{C}_{t}$, which is pure geometry, and we’ll use the range of $V_{t}$-norm to determine the range of $V_{t}^{-1}$-norm.
-Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$.
-(a): $\mathcal{C}_{t}$ is the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
-(b): As the decision-maker collects more and more data, the confidence ellipsoid $\mathcal{E}_{t}$ shrinks accordingly (here different color corresponds to different levels of exploration).
-The potentially optimal actions point to similar directions, which essentially determine their $V_{t}$-norm.
-Hence it’s plausible one could get a proper range of their $V_{t}$-norm using geometry.
-The following main theorem of this paper derives an upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Section 5.3.
+Nevertheless, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$.
+Therefore, it is possible to estimate the range of the $V_{t}$-norm by employing geometric reasoning.
+Subsequently, this estimated range will be utilized to ascertain the range of the $V_{t}^{-1}$-norm.
+Figure 2: Illustration of potentially optimal actions set $\mathcal{C}_{t}$ in $\mathbb{R}^{2}$.
+(a): $\mathcal{C}_{t}$ is $\mathcal{E}_{t}$’s projection onto ${\mathcal{S}}_{d-1}$.
+(b): As more data is collected, $\mathcal{E}_{t}$ shrinks (colors show exploration levels).
+Potentially optimal actions point in similar directions, determining their $V_{t}$-norm.
+This suggests their $V_{t}$-norm range could be estimated geometrically.
+The main theorem (proved in Section 9) derives an upper bound for $\alpha_{t}$ based on this idea.
 Theorem 2
-Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$. Define $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$ and
-[…]
-[…]
-where we let
-$\quad m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t}/\lambda_{d}(V_{t}))^{2},\quad M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t})^{2})$
-and $k\in[d]$ is such that $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$.
-Then for all $t\in[T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$, it holds for all $s\leq t$ that
-$\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s},\quad\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s},\quad\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}.$
-To better understand what Theorem 2 implies, we discuss some special cases.
-Case 1: a pure exploration regime.
-When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$.
-Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$.
-Specifically, if $V_{t}=D\mathbb{I}_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
-Case 2: linear structures.
-When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated in Appendix 11).
-To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small.
-The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1.
-This provides a plausible explanation for the empirical success of LinTS and Greedy.
-Proposition 5
-Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{\max}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
+Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$.
+Define $m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}/\lambda_{d}(V_{t}))^{2}$, $M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}})^{2})$.
+Let $k\in[d]$ be the integer that satisfies $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$.
+Define $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and
+$\Phi_{t}$ =$=\begin{cases}(\lambda_{1}^{-1}(V_{t})+\lambda_{d}^{-1}(V_{t})-m_{t}\lambda_{1}^{-1}(V_{t})\lambda_{d}^{-1}(V_{t}))^{\frac{1}{2}},&\ \ \ \ \textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}\\ \lambda_{d}^{-\frac{1}{2}}(V_{t}),&\ \ \ \ \textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}<\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}\end{cases},$
+$\Psi_{t}$ =$=\begin{cases}(\lambda_{k}^{-1}(V_{t})+\lambda_{k+1}^{-1}(V_{t})-M_{t}\lambda_{k}^{-1}(V_{t})\lambda_{k+1}^{-1}(V_{t}))^{\frac{1}{2}},&\textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\\ \lambda_{1}^{-\frac{1}{2}}(V_{t}),&\textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}<\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\end{cases}.$
+Then for all $t\in[T]$, conditioned on $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$, it holds for all $s\leq t$ that $\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}$.
+To better understand what Theorem 2 implies, we discuss some special cases in Appendix 14 and provide empirical validations for them.
+6 A Meta-Algorithm for Course-Correction
+This section demonstrates how the data-driven regret bound can enhance standard bandit algorithms.
+We propose a meta-algorithm that creates course-corrected variants of base algorithms, achieving minimax-optimal frequentist regret guarantees while preserving most original characteristics, including computational efficiency and typically low regret.
+We take LinTS as an example of the base algorithm, and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR).
+The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions.
+Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$.
+If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action.
+We remark that setting $\iota_{t}=0$ for all $t\in[T]$ yields the corresponding Greedy-MR algorithm.
+The pseudocode is presented in Algorithm 2 in Appendix 15.
 Remark 3
-Clearly, one can further improve the bound in Equation (6) by plugging in the selected action $\widetilde{x}_{t}$.
-This is true for standard algorithms such as OFUL, LinTS, and Greedy.
-Explicitly, the uncertainty ratio can be bounded using $\widetilde{x}_{t}$ and Theorem 2 as
-$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}\leq\frac{\Phi_{t}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}.$ (7)
-However, for the variants of POFUL that we are going to propose in the Section 6, it is necessary to calculate the bound before selecting the action $\widetilde{x}_{t}$.
-Therefore, the bound shouldn’t depend on $\widetilde{x}_{t}$.
-5.3 Sketch of the Proof
+Computing $\widehat{\mu}_{t}$ primarily requires SVD decomposition of the sample covariance matrix $V_{t}$.
+Since $V_{t}=\lambda_{\text{reg}}I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ is updated via rank-one matrices, its SVD can be efficiently updated (Gandhi and Rajgor, 2017), preventing computational bottlenecks.
+By design, course-corrected algorithms ensure that $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$.
+Substituting this upper bound into Theorem 1 establishes that these algorithms achieve optimal frequentist regret, up to a constant factor.
+Corollary 1
+TS-MR and Greedy-MR achieve a frequentist regret of $\widetilde{\mathcal{O}}(\max\{\mu,2\}d\sqrt{T})$.
+In high-risk settings where LinTS and Greedy may fail, a small $\mu$ ensures TS-MR and Greedy-MR select more OFUL actions, promoting sufficient exploration.
+Conversely, in low-risk settings where the original algorithms perform well, a large $\mu$ favors TS and greedy actions, minimizing unnecessary exploration and reducing computational cost.
+In Appendix 16, we show how $\mu$ impacts the fraction of OFUL actions in TS-MR and Greedy-MR and their performance.
+Results indicate that course-corrected algorithms maintain robustness across a range of moderate $\mu$ values, suggesting that the precise selection of $\mu$ is unlikely to present a significant practical concern.
+7 Simulations
+We aim to compare TS-MR, Greedy-MR, and key baseline algorithms, via simulation.
+7.1 Synthetic datasets
+We conduct simulations on three representative synthetic examples. We average simulation results over 100 independent runs for each of the examples. The results are shown in Figure 3.
+Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
+We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm.
+At each time $t$, we generate 100 i.i.d.
+random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$.
+This is a basic example of standard stochastic linear bandit problems without any extra structure.
+We set the threshold $\mu=8$ for TS-MR and Greedy-MR.
+TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
+Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
+We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$.
+At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,I_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0.
+Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks.
+This problem is equivalent to a 10-armed contextual bandit.
+We set $\mu=12$ for TS-MR and Greedy-MR.
+In this setting, Greedy performs suboptimally due to a lack of exploration for some arms.
+Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects large regret proxy $\widehat{\mu}_{t}$.
+Example 3. Prior mean mismatch (Hamidi and Bayati, 2020b).
+This is an example in which LinTS is shown to incur linear Bayesian regret.
+We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},I_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}e_{i}/\sqrt{3d},\quad x_{b}=\sum_{i=11}^{3d}e_{i}\sqrt{3d}-\sum_{i=1}^{d}e_{i}/\sqrt{3d}$.
+It is shown in Hamidi and Bayati (2020b) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$.
+Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action.
+We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit.
+We set $\mu=12$ for TS-MR and Greedy-MR.
+We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
+(a) example 1
+(b) example 2
+(c) example 3
+(d) example 1
+(e) example 2
+(f) example 3
+Figure 3: Simulation results on synthetic data.
+(a) - (c): Cumulative regret of TS-MR and Greedy-MR versus baseline algorithms.
+Shaded regions show ±$\pm 2$ SE of mean regret.
+(d) - (f): Fraction of OFUL actions in TS-MR and Greedy-MR.
+7.2 Real-world datasets
+We explore the performance of standard POFUL algorithms and the proposed TS-MR and Greedy-MR algorithms on real-world datasets.
+We use three classification datasets from OpenML: Cardiotocography, JapaneseVowels, and Segment, representing healthcare, pattern recognition, and computer vision domains.
+Following Bietti et al.
+(2021); Bastani et al.
+(2021), we convert these classification tasks to contextual bandit problems and embed them into linear bandit problems as in Example 2, Section 7.1.
+Each class becomes an action where the decision-maker receives a binary reward (1 for correct classification, 0 otherwise) plus Gaussian noise.
+We plot the cumulative regret (averaged over 100 runs) for all algorithms.
+Figure 4 shows that for all real-world datasets: OFUL and TS-Freq perform poorly due to their conservative exploration; LinTS and Greedy are achieving empirical success even though they don’t have theoretical guarantees; TS-MR and Greedy-MR retain the desirable empirical performance of LinTS and Greedy, while enjoying the minimax optimal frequentist regret bound.
+(a) Cardiotocography dataset
+(b) JapaneseVowels dataset
+(c) Segment dataset
+(d) Cardiotocography dataset
+(e) JapaneseVowels dataset
+(f) Segment dataset
+Figure 4: Simulation results on real-world datasets.
+(a) - (c): Cumulative regret of all algorithms.
+Shaded regions show the ±$\pm 2$ SE of the mean regret.
+(d) - (f): Fraction of OFUL actions of TS-MR and Greedy-MR.
+Remark 4
+Simulation results in Figures 3 and 4 show OFUL actions are primarily used in the early stages.
+This indicates: (1) Greedy-MR and TS-MR implement OFUL actions only when necessary, maintaining a low OFUL fraction throughout most of the time horizon, substantially reducing computational cost; and (2) limited course-corrected exploration at the beginning efficiently remedies TS and Greedy in problematic instances.
+8 Conclusion
+In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
+Our approach allows for the computation of a data-driven frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm.
+Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest.
+As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail.
+We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
+{APPENDICES}
+9 Sketch of the Proof
 In this section, we present the proof of Theorem 2.
-In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
+In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}\right\}$ to represent $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and $\mathcal{E}^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, since the proof works for both of them.
 First note that, when $\|\widehat{\theta}_{t}\|_{V_{t}}<\beta_{t}$, the bound in Theorem 2 becomes
@@
 By inserting the upper and lower bounds of the $V_{t}$-norm from Lemmas 1 and 2 into Lemma 3, we finalize the proof of Theorem 2.
-6 A Meta-Algorithm for Course-Correction
-In this section, we show how the data-driven regret bound can be utilized to improve standard bandit algorithms.
-Specifically, we propose a meta-algorithm that constructs course-corrected variants of the base algorithms.
-This ensures they achieve a minimax-optimal frequentist regret guarantee (up to constant factors) while retaining the majority of the original characteristics of the base algorithms, such as computational efficiency and low regret in most cases.
-We take LinTS as an example and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR).
-The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions when necessary.
-Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$.
-If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action.
-We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm.
-The pseudocode is presented in Algorithm 2.
-Algorithm 2 TS-MR (Greedy-MR)
-0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
-  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
-  for t = 1, 2 , …, T  do
-     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
-     $\widehat{\mu_{t}}\leftarrow\widehat{\alpha}_{t}(1+\iota_{t})+1+\iota_{t}$
-     if $\widehat{\mu}_{t}\leq\mu$ then
-        Sample $\eta_{t}\sim\mathcal{D}^{SA}(\delta^{\prime})$ (defined in Section 3)
-        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$
-        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
-     else
-        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})}\langle{x},{\widetilde{\theta}_{t}}\rangle$
-     end if
-     Observe reward $r_{t}$
-     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
-     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
-  end for
-By construction, the course-corrected algorithms have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$.
-Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
-Corollary 1
-TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
-Remark 4
-In practice, the choice of the threshold $\mu$ depends on the problem settings.
-In a high-risk setting where LinTS and Greedy fail with high probability, one can set a small $\mu$ so that TS-MR and Greedy-MR select more OFUL actions to guarantee the necessary amount of exploration.
-In a low-risk setting where original LinTS and Greedy algorithms work well, one can set a large $\mu$ and hence TS-MR and Greedy-MR select TS and greedy actions respectively to avoid unnecessary exploration and save the computational cost.
-7 Simulations
-In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
-7.1 Synthetic datasets
-(a) example 1
-(b) example 2
-(c) example 3
-Figure 3: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
-Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
-We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm.
-At each time $t$, we generate 100 i.i.d.
-random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$.
-This is a basic example of standard stochastic linear bandit problems without any extra structure.
-We set the threshold $\mu=8$ for TS-MR and Greedy-MR.
-We average simulation results over 100 independent runs.
-The results are summarized in Figure 2(a).
-TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
-Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
-We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$.
-At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0.
-Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks.
-This problem is equivalent to a 10-armed contextual bandit.
-We ser $\mu=12$ for TS-MR and Greedy-MR.
-Again, we average simulation results over 100 independent runs.
-The results are summarized in Figure 2(b).
-In this setting, Greedy performs suboptimally due to a lack of exploration for some arms.
-Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects large regret proxy $\widehat{\mu}_{t}$.
-Example 3. Prior mean mismatch (Hamidi and Bayati, 2020b).
-This is an example in which LinTS is shown to incur linear Bayesian regret.
-We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$.
-It is shown in Hamidi and Bayati (2020b) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$.
-Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action.
-We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit.
-We ser $\mu=12$ for TS-MR and Greedy-MR.
-We carried out 100 independent runs.
-The results are summarized in Figure 2(c).
-We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
-7.2 Real-world datasets
-We explore the performance of standard POFUL algorithms and the proposed TS-MR and Greedy-MR algorithms on real-world datasets.
-We consider three datasets for classification tasks on the OpenML platform, Cardiotocography, JapaneseVowels and Segment.
-They focus on healthcare, pattern recognition, and computer vision problems respectively.
-Setup.
-We follow the standard approach in the literature that converts classification tasks to contextual bandit problems, and then embeds it into linear bandit problems in the same way as Example 2 in Section 6.
-Specifically, we regard each class as an action so that at each time, the decision-maker assigns a class to the feature observed and receives a binary reward, namely 1 for assigning the correct class and 0 otherwise, plus a Gaussian noise.
-We plot the cumulative regret (averaged over 100 duplicates) for all algorithms.
-Figure 4 shows that for all real-world datasets: OFUL and TS-Freq perform poorly due to their conservative exploration; LinTS and Greedy are achieving empirical success even though they don’t have theoretical guarantees; TS-MR and Greedy-MR retain the desirable empirical performance of LinTS and Greedy, while enjoying the minimax optimal frequentist regret bound.
-(a) Cardiotocography dataset.
-(b) JapaneseVowels dataset.
-(c) Segment dataset.
-Figure 4: Cumulative regret of all algorithms on real-world datasets.
-8 Conclusion
-In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
-Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm.
-Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest.
-As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail.
-We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
-{APPENDICES}
 10 Proof of Lemmas for Theorem 2
@@
 By the spectral theorem, $V$ can be decomposed as $V=U\Lambda U^{\top}$, where the columns of $U$ consist of orthonormal eigenvectors of $V$, denoted by $\{u_{11},\dots,u_{1n_{1}},u_{21},\dots,u_{2n_{2}},\dots,u_{N_{V}1},\dots,u_{N_{V}n_{N_{V}}}\}$, where $n_{1},\dots,n_{N_{V}}$ are the algebraic multiplicity of the eigenvalues respectively.
-Since $V$ is a symmetric matrix, the eigenvectors for a basis of $\mathbb{R}^{d}$ and we have $\sum_{i=1}^{N_{V}}n_{i}=d$.
+Since $V$ is a symmetric matrix, the eigenvectors form a basis of $\mathbb{R}^{d}$ and we have $\sum_{i=1}^{N_{V}}n_{i}=d$.
 We can write $x$ as a linear combination $x=\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}u_{ij}$, where $\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}^{2}=1$.
@@
 The KKT conditions are given by
-$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{N_{V}}a_{i}=1,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$ (10)
+$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\forall i,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{N_{V}}a_{i}=1,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$ (9)
 To satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices.
 Hence, $a_{i}$ can only be non-zero for at most two distinct eigenvalues, denoted by $\lambda_{i}$ and $\lambda_{j}$, where $i<j$ and $\lambda_{i}>\lambda_{j}$.
-Namely, the solution to (10) lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{i}$ and $\lambda_{j}$.
+Namely, the solution to (LABEL:eq:KKT) lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{i}$ and $\lambda_{j}$.
 Let $y=\|x\|_{V}^{2}$, we have $a_{i}\lambda_{i}+a_{j}\lambda_{j}=y$ and $a_{i}\lambda_{i}^{-1}+a_{j}\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
@@
 Proof.
-By Propostion 1, we have
-${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
-$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
+By Proposition 1, we have
+${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\geq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right\}\right]$
+$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\geq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right]$
 $\geq 1-\frac{\delta}{2}.$
 Similarly, by Definition 2, we have
-${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
-$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
+${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\notin\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}\right\}\right]$
+$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\notin\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}\right]$
 $\geq 1-\frac{\delta}{2}.$
@@
 Proof.
-Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
-$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (11)
-In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. We have
+Since ${\theta}^{\star}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$, it holds that
+$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS},\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}=\iota_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$ (10)
+We have
 $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $=\left(\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle\right)+\left(\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\right)$
@@
 To bound the second term on the right hand side, recall by Equation (3) the POFUL action $\widetilde{x}_{t}$ satisfies
-$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
-Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
-The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (11). As a result, we have
-$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
-$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t},$
-$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
-Combing all terms above, we have
-$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
+$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
+Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$.
+The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle inequality of the $V_{t}^{-1}$-norm, and the concentration condition (10). As a result, we have
+$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS},$
+$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}},$
+$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}.$
+Combining all terms above, we have
+$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}.$
 11.3 Proof of Theorem 1
@@
 $\mathcal{R}(T)$ $\leq\sum_{t=1}^{T}\left(\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right)\mathbb{I}\{\mathcal{A}_{t}\}$
-$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
+$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
 Recall $\|x^{\star}_{t}\|_{V^{-1}_{t}}=\alpha_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}$ and $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$, we have
-$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
-Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
-$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
+$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
+Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}=\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$, we obtain
+$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
 This concludes the proof. ∎
-10.4 Proof of Proposition 5
+12 Regret Decomposition of POFUL
+To discuss the relationship between our method and those based on optimism (Abbasi-Yadkori et al., 2011; Agrawal and Goyal, 2013; Abeille et al., 2017), we decompose the regret of POFUL into three terms, and sketch how they are bounded separately.
+Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the optimal action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
+$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
+$=\underbrace{\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle}_{\mathcal{R}^{PE}(T)}+\underbrace{\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle}_{\mathcal{R}^{PVT}(T)}+\underbrace{\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle}_{\mathcal{R}^{RLS}(T)}\,.$
+$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator.
+Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate.
+By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda_{\text{reg}})$.
+Hence, the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which, via the Cauchy-Schwarz inequality, is characterized by the radius of the confidence ellipsoid $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$.
+$\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}$ rather than using $\widehat{\theta}_{t}$ as OFUL would.
+Similar to $\mathcal{R}^{RLS}(T)$, the upper bound for this term is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$.
+As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$.
+Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound.
+For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$.
+However, it is known that, see for example (Agrawal and Goyal, 2013; Abeille et al., 2017; Hamidi and Bayati, 2020b), such selection could be problematic for bounding $\mathcal{R}^{PE}(T)$.
+$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term.
+In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution, which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$.
+In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$.
+For OFUL, this term is bounded properly with high probability by the optimistic selection of OFUL actions [^1] The term is further split into two parts as shown in the first few lines of proof of Proposition 4 in Appendix 11.2, with one term bounded by the optimism and the other term controlled by the $V_{t}$ distance between $\theta^{\star}$ and $\widetilde{\theta}_{t}$..
+For LinTS, the only known analysis due to Agrawal and Goyal (2013) and Abeille et al.
+(2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$.
+The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ to enforce exploration.
+For example, Abeille et al.
+(2017) demonstrate that by using inflation, LinTS samples optimistic $\widetilde{\theta}_{t}$ with a probability greater than a constant, which means that the inequality $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ holds with a constant probability.
+For Greedy, there is no general theoretical guarantee for bounding the pessimism term without imposing additional structural assumptions.
+Our approach to tackle this challenge distinguishes us from methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017).
+Interestingly, Abeille et al.
+(2017) conjectured that non-optimistic samples may provide beneficial exploration and help control the growth of regret.
+By relaxing the requirement for optimistic samples, we avoid inflating the posterior, which in turn prevents a $\sqrt{d}$-gap in the regret.
+Instead, we introduce a measure of the “quality” of POFUL actions with respect to the regret and develop a computable upper bound for this quality using geometric information in the data.
+This approach yields a data-driven regret bound for POFUL that matches the minimax optimal regret in some settings.
+Furthermore, it allows for the construction of variants of LinTS and Greedy with a provable frequentist regret bound that is minimax optimal up to only a constant factor.
+13 Discrete Action Space
+This section introduces the method for establishing an upper bound on $\alpha_{t}$ in the context of discrete action spaces.
+It turns out that the process involves comparing confidence interval lengths across all potentially optimal actions.
+To see this, recall for an arbitrary action $x\in\mathcal{X}_{t}$, by Proposition 1, with high probability its expected reward is bounded as
+$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\eqqcolon U_{t}(x).$
+The confidence interval length is proportional to $\|x\|_{V_{t}^{-1}}$.
+Therefore, to characterize the range of $\|x_{t}^{\star}\|_{V_{t}^{-1}}$, we only need to identify the set of potentially optimal actions and calculate the minimal $V_{t}^{-1}$-norm among all actions in that set.
+Note that for an action $x\in\mathcal{X}_{t}$ to remain potentially optimal, it cannot be dominated by another action.
+I.e., we require $U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}L_{t}(y)$.
+Therefore, define
+$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\},$
+the set of potentially optimal actions is given by $C_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS})$.
+Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT})$, where we let $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$.
+The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
+By comparing the range of $V_{t}^{-1}$-norm for both sets, an upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$ is given by
+$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS})}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT})}\|x\|_{V_{t}^{-1}}}.$ (11)
+14 Example Cases and empirical validations of Theorem 2
+To better understand what Theorem 2 implies, we discuss some special cases in this section.
+Case 1: a pure exploration regime.
+When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$.
+Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$.
+Specifically, if $V_{t}=DI_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
+Case 2: linear structures.
+When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated below).
+To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small.
+The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1.
+This provides a plausible explanation for the empirical success of LinTS and Greedy.
+Proposition 5
+Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{1}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
 Proof.
-This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3.
-Without loss of generality, we let $M>\lambda_{2}$.
-Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$.
+This corresponds to the case where $m,M\to\lambda_{1}(V_{t})$ in Lemma 3.
+Without loss of generality, we let $M>\lambda_{2}(V_{t})$.
+Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}(V_{t})+\lambda_{d}^{-1}(V_{t})-m{\lambda_{1}^{-1}(V_{t})\lambda_{d}^{-1}(V_{t})})/(\lambda_{k}^{-1}(V_{t})+\lambda_{k+1}^{-1}(V_{t})-{M}{\lambda_{k}^{-1}(V_{t})\lambda_{k+1}^{-1}(V_{t})})\to\lambda_{1}^{-1}(V_{t})/\lambda_{1}^{-1}(V_{t})=1$.
 ∎
-11 Empirical Validation of Case 2 from Section 5.2
-To empirically confirm that Case 2, as discussed in Section 5.2, is not merely theoretical, we examine the condition outlined in Proposition 5.
+To empirically confirm that Case 2 is not merely theoretical, we examine the condition outlined in Proposition 5.
 Our focus is on the ratio $\zeta_{t}\coloneqq\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}/\|\widehat{\theta}_{t}\|^{2}}{\lambda_{1}(V_{t})}$.
@@
 Specifically, a $\zeta_{t}$ value of 1 signifies that $\widehat{\theta}_{t}$ is within the top eigenspace.
-We re-evaluate Example 1 from Section 7, which epitomizes the general linear bandit problem in its standard form. The empirical sequence ${\zeta_{t}}$ is depicted in Figure 5.
+We resort to Example 1 from Section 7, which epitomizes the general linear bandit problem in its standard form. The empirical sequence ${\zeta_{t}}$ is depicted in Figure 6.
 Figure 6: Evolution of the alignment proxy $\zeta_{t}$ in Example 1.
@@
 It is important to note that the initial sharp decline observed is attributed to the behavior of the regularized least squares estimator in over-parameterized scenarios, when $t$ is less than the dimension $d$.
+15 TS-MR and greedy-MR Algorithms
+Algorithm 2 TS-MR (Greedy-MR)
+0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
+  Initialize $V_{0}\leftarrow\lambda I_{d}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
+  for t = 1, 2 , …, T  do
+     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
+     $\widehat{\mu_{t}}\leftarrow\widehat{\alpha}_{t}(1+\iota_{t})+1+\iota_{t}$
+     if $\widehat{\mu}_{t}\leq\mu$ then
+        Sample $\eta_{t}\sim\mathcal{D}^{SA}(\delta^{\prime})$ (defined in Section 3)
+        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-\frac{1}{2}}\eta_{t}$
+        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
+     else
+        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
+     end if
+     Observe reward $r_{t}$
+     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
+     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
+  end for
+16 Influence of $\mu$
+In this section, we investigate the influence of $\mu$ and discuss strategies for selecting its value.
+We vary the value of $\mu$ and conduct simulations on Example 2 and the Segment dataset as done in Section 7.
+The former represents a scenario where course-correction is necessary (otherwise LinTS fails), while the latter represents a practical scenario where course-correction is unnecessary.
+The results are shown in Figures 7 and 8.
+(a) example 2, $\mu=4$
+(b) example 2, $\mu=8$
+(c) example 2, $\mu=12$
+(d) example 2, $\mu=4$
+(e) example 2, $\mu=8$
+(f) example 2, $\mu=12$
+Figure 7: Cumulative regret and fraction of OFUL actions of TS-MR and Greedy-MR on example 2. Shaded regions show the ±$\pm 2$ SE of the mean regret.
+(a) Segment dataset, $\mu=4$
+(b) Segment dataset, $\mu=8$
+(c) Segment dataset, $\mu=12$
+(d) Segment dataset, $\mu=4$
+(e) Segment dataset, $\mu=8$
+(f) Segment dataset, $\mu=12$
+Figure 8: Cumulative regret and fraction of OFUL actions of TS-MR and Greedy-MR on Segment datasets. Shaded regions show the ±$\pm 2$ SE of the mean regret.
+We have the following observations on the influence of $\mu$:
+  • In general, as $\mu$ increases, the fraction of OFUL actions decreases.
+For a sufficiently small $\mu$, the algorithms become OFUL (e.g., TS-MR in Figures 7(a) and 8(a)).
+For a large $\mu$, the algorithms are similar to the original algorithms (e.g., TS-MR and Greedy-MR in Figures 7(c) and 8(c)).
+  • By optimizing the choice of $\mu$, the performance of the corrected algorithm might outperform both OFUL and the base algorithm (e.g., Greedy-MR in Figures 7(b) and 8(b)).
+This demonstrates that TS/Greedy-MR are not merely naive interpolations between TS/Greedy and OFUL, and that proper exploration in the initial stage benefits the long-term performance of the algorithms.
+  • As long as $\mu$ isn’t too small, the performance of the course-corrected algorithm are relatively robust to the choice of $\mu$. The simulations show $\mu\in[8,12]$ is a good initial choice.
+Empirical guidelines and heuristics for choosing $\mu$.
+In practice, we recommend setting $\mu$ to a moderate value, typically within the range $[8,12]$, as validated by our simulation results.
+The selection of $\mu$ should be guided by heuristics that ensure an appropriate fraction of course-corrected exploration during the initial stage.
+A suitable $\mu$ value is indicated when the fraction of OFUL actions is high at the beginning and gradually decreases to and maintains a low level.
+If the fraction of OFUL actions remains consistently high, increasing $\mu$ can help reduce computational costs.
+Conversely, if very few OFUL actions are executed during the initial stage, decreasing $\mu$ may be beneficial.
+Finally, we note that since algorithmic performance remains robust across moderate values of $\mu$, the precise selection of $\mu$ is unlikely to be a significant practical concern.